### v1

# On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral

Abstract
Tool-integrated (TI) reinforcement learning (RL) enables large language models (LLMs) to perform multi-step reasoning by interacting with external tools such as search engines and retrievers. Group Relative Policy Optimization (GRPO), exemplified by the recent Search-R1 (jin2025search), offers fast convergence and a value-free formulation that makes it appealing for this setting, yet consistently suffers from training collapse. We identify Lazy Likelihood Displacement (LLD), a systematic reduction or stagnation in the likelihood of both correct and incorrect responses, as the core mechanism driving this failure. LLD emerges early and triggers a self-reinforcing LLD Death Spiral, where declining likelihood leads to low-confidence responses, inflating gradients and ultimately causing collapse. We empirically characterize this process across models on a Search-R1-style, search-integrated question answering task, revealing a consistent three-phase trajectory: early stagnation, steady decay, and accelerated collapse. To address this, we propose a lightweight likelihood-preserving regularization $\rm LLDS$ for GRPO that activates only when a trajectory’s likelihood decreases, and regularize only the tokens responsible. This fine-grained structure mitigates LLD with minimal interference to optimization. Across seven open-domain and multi-hop QA benchmarks, our method stabilizes training, prevents gradient explosion, and yields substantial performance improvements, including +37.8% gains on Qwen2.5-3B and +32.0% gains on Qwen2.5-7B. Our results establish LLD as a fundamental bottleneck in GRPO-based TIRL and provide a practical path toward stable, scalable training of tool-integrated LLM.
1 Introduction
Figure 1: Comparative performance of LLDS and baseline methods on benchmark datasets. All baselines are built upon Qwen2.5-7B-Instruct. See Sec.˜6.1 for details.
Large language models (LLMs) increasingly leverage external tools, such as search engines (jin2025search) and code-execution environments (feng2025retool; zeng2025simplerl), to augment their reasoning capabilities. This tool-integrated reasoning (TIR) paradigm has driven recent progress across factual question answering (jin2025search), image-based reasoning (jiang2025verltool), and mathematical problem solving (feng2025retool; jiang2025verltool). By enabling models to iteratively query tools, verify intermediate steps, and refine their thoughts, tool calls substantially elevate reasoning quality (qin2023toolllm). These advances naturally motivate the use of reinforcement learning (RL) to train LLMs to plan, interact with tools, and master multi-step decision making, as instantiated by recent GRPO-based frameworks such as Search-R1 (jin2025search).
However, extending TIR to the RL setting introduces severe and persistent training instability. In particular, GRPO training in Search-R1-style jin2025search, search-integrated QA pipelines frequently exhibit sudden reward drops and catastrophic collapse (jin2025search; sun2025zerosearch). These failures are especially pronounced in multi-turn settings (xue2025simpletir), where tool feedback becomes part of the model’s conditioning context, creating long, highly entangled trajectories. Although prior work has observed these failures, the underlying mechanisms remain poorly understood.
In this work, we firstly identify Lazy Likelihood Displacement (LLD) (deng2025effect; razin2024unintentional; ren2024learning), the stagnation or reduction of likelihood for both correct and incorrect responses during GRPO optimization (deng2025effect), as the fundamental but previously overlooked source of collapse in Search-R1–style tool-integrated reinforcement learning (TIRL). Using search-integrated question answering as a case study (see Fig.˜2), we show that LLD emerges early and persistently: even as rewards increase, the likelihood of correct responses enters a monotonic decline. This behavior appears across model scales, indicating that LLD is a structural failure mode of GRPO-based TIRL rather than a configuration-specific artifact. We further demonstrate that LLD drives a self-reinforcing LLD death spiral, where the resulting low-confidence regime amplifies negative-gradient influence from incorrect trajectories, accelerating likelihood decay, triggering entropy spikes, inflating likelihood ratios, and ultimately causing the large-gradient instability that leads to collapse.
To counteract this failure mode, we propose a simple yet effective likelihood-preserving regularization $\rm LLDS$ that prevents harmful likelihood reductions. Our method integrates seamlessly with GRPO and introduces two layers of selectivity: (i) response-level gating, which activates the regularization only when a trajectory’s overall likelihood decreases, and (ii) token-level selectivity, which penalizes only the tokens responsible for the decrease. This fine-grained design directly mitigates LLD while minimally interfering with GRPO’s optimization behavior. By preventing unintentional downward likelihood drift, $\rm LLDS$ keeps optimization away from the unstable high-gradient regime and maintains healthy training dynamics. Empirically, this translates into stable training, suppressed gradient explosion, and consistent gains across seven open-domain and multi-hop QA benchmarks. Our main contributions are threefold:
  • We conduct the first systematic investigation of LLD in GRPO-driven TIRL and show that it arises with striking frequency, consistently progressing through a characteristic failure trajectory: steady likelihood decay, followed by gradient amplification and entropy explosion, and ultimately collapse.
  • We propose a lightweight likelihood-preserving regularization $\rm LLDS$ that selectively regularize likelihood reductions and resolved the collapse issue in GRPO training.
  • By stabilizing training, we realize significant performance gains across seven QA benchmarks (see Fig.˜1), demonstrating a more robust and reliable approach to RL-driven TIR.
Figure 2: We illustrate the likelihood displacement in tool-integrated RL training. The steady-decay phase (60-120) emerges when the reward begins to increase only gradually. In the subsequent acceleration phase (after step 120), the likelihood of correct responses drops sharply, accompanied by a sudden surge in gradient magnitude (red star), leading to gradient explosion. A zoomed-in view of the acceleration region further highlights this effect, showing a clearer likelihood displacement, where the gradient accelerates rapidly while the reward starts to decline.
2 Related Work
Tool-Integrated Reasoning and Agentic LLMs. Tool use has emerged as a powerful paradigm for equipping LLMs with adaptive reasoning capabilities. Early approaches relied on prompt-based orchestration (lu2023chameleon; shen2023hugginggpt) or multi-agent delegation frameworks to invoke tools without explicit training. Instruction-tuned models (gou2023tora; qin2023toolllm) later introduced structured tool-calling behaviors via supervised learning, but these systems remained largely static and were constrained to single-turn interactions. More recent work demonstrates that reinforcement learning can substantially enhance tool integration by enabling models to learn tool-usage policies through environmental feedback and task success. Notable systems such as RETool (feng2025retool), VERL-Tool (jiang2025verltool), and agentic LLM frameworks (maiagentic) support multi-step reasoning through dynamic tool use, self-verification, and error correction. This transition from static instruction-following to feedback-driven optimization has proven effective across a range of domains, including mathematical problem solving with code execution (xue2025simpletir), open-domain question answering with retrieval (jin2025search), SQL generation from natural language (jiang2025verltool), and multi-modal visual reasoning (gao2024multi).
Training Collapse in Tool-Integrated GRPO. GRPO (guo2025deepseek) has gained popularity in reinforcement learning due to its value-free, outcome-driven formulation. However, applying GRPO to train LLMs for multi-turn tool-integrated reasoning remains highly unstable (jin2025search). When initialized from base models using only verifiable rewards (Zero RL), GRPO-trained models often experience catastrophic failure, marked by abrupt reward collapse (xue2025simpletir; maiagentic; jin2025search). In contrast, Proximal Policy Optimization (PPO) (schulman2017proximal) tends to exhibit more stable behavior under comparable settings (jin2025search). These collapse dynamics have been consistently observed across various systems, including Search-R1 (jin2025search), SimpleTIR (xue2025simpletir), and ZeroSearch (sun2025zerosearch). While prior work has largely focused on empirical observations of failure, only xue2025simpletir attempt to explain the phenomenon, attributing it to low-likelihood incorrect responses, which inflate importance weights and destabilize learning. Yet this explanation does not fully account for the structural causes of such low-likelihood trajectories. Notably, we observe that reward degradation often precedes the final collapse, suggesting deeper optimization issues. In this work, we identify Lazy Likelihood Displacement (LLD) as the core mechanism underlying GRPO’s failure mode in multi-turn TIRL.
3 Priliminary
TIRL enables an LLM to interact with the external tools. We denote the query by $\bm{x}$, the tool feedback by $\bm{o}$, and the model’s action by $\bm{y}$. For a given query $\bm{x}$, the model autoregressively generates an action at turn $t$ according to
$\pi_{\theta}\!\left(\bm{y}_{t}\mid\bm{x},\bm{y}_{0},\bm{o}_{0},\bm{y}_{1},\ldots,\bm{o}_{t-1}\right),$
where $\bm{y}_{t}$ is produced based on the accumulated context, and $\bm{o}_{t-1}\sim\mathcal{T}$ denotes the tool feedback returned at turn t1t-1 if the previous action $\bm{y}_{t-1}$ invoked a tool $\mathcal{T}$ (e.g., issuing a search query). If no tool is invoked, $\bm{o}_{t-1}$ is an empty string. The multi-turn setting therefore corresponds to trajectories that involve multiple tool calls and, consequently, multiple rounds of tool feedback. However, tool feedback $\mathbf{o}_{t}$ is inherently out-of-distribution (OOD) for the pretrained language model, since such feedback originates from external environments rather than the model’s own generative distribution. Directly forcing the model to fit these feedback tokens introduces a large training–inference mismatch and may cause the model to memorize environment-specific feedback. Therefore, previous works commonly mask out feedback tokens during supervised training to stabilize optimization and avoid overfitting to external content (jin2025search).
Tool-Integrated GRPO with Feedback Mask.
The GRPO loss, introduced in DeepSeek-Math (shao2024deepseekmath) and DeepSeek-R1 (guo2025deepseek), enhances policy optimization by refining how reward and loss signals are computed. Unlike Proximal Policy Optimization (PPO) (schulman2017proximal), GRPO removes the need for value function estimation and instead uses group-relative reward normalization for optimization. Specifically, for a query–answer pair $(\bm{x},\bm{a})$, the policy $\pi_{\theta}$ samples $G$ responses:
$\left\{(\bm{y}_{i,0},\bm{o}_{i,0},\ldots,\bm{y}_{i,t},\bm{o}_{i,t},\ldots,\bm{o}_{i,T_{i}-1},\bm{y}_{i,T_{i}})\right\}_{i=1}^{G},$
where $T_{i}$ denotes the number of tool calls, and thus determines the number of multi-turn interactions in the $i$-th rollout. Each action $\bm{y}_{i,t}$ consists of $|\bm{y}_{i,t}|$ tokens, and we denote by $\bm{y}_{i,t,<k}$ the prefix consisting of its first k1k-1 tokens. Let $r_{i}$ denote the reward assigned to the $i$-th response. The advantage for the $t$-th action of $i$-th response is defined via group-level normalization:
$\hat{A}_{i,t,k}:=\frac{r_{i}-\mu}{\sigma},\qquad k=1,\ldots,|\hat{\bm{y}}_{i,t}|,$
where $\mu=\widehat{\mathbb{E}}[\{r_{i}\}_{i=1}^{G}]$ and $\sigma=\sqrt{\widehat{\mathrm{Var}}[\{r_{i}\}_{i=1}^{G}]}$ are the empirical mean and standard deviation of rewards within the group. Each token of the same trajectory shares the same normalized advantage. The tool-integrated GRPO objective with feedback mask is given by:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}(\bm{x},\bm{a})\sim\mathcal{D}\\ \{\bm{y}_{i,0},\ldots,\bm{y}_{i,t},\bm{o}_{i,t},\bm{y}_{i,T_{i}}\}\sim(\pi_{\theta_{\mathrm{old}},\mathcal{T}})\end{subarray}}$
$\Bigg[\frac{1}{\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}|\hat{\bm{y}}_{i,t}|}\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}\sum_{k=1}^{|\hat{\bm{y}}_{i,t}|}\min\!\left(\gamma_{i,t,k}(\theta)\,\hat{A}_{i,t,k},\;\hat{A}_{i,t,k}\,\mathrm{clip}\big(\gamma_{i,t,k}(\theta),\,1-\varepsilon,\,1+\varepsilon\big)\right)\Bigg],$ (1)
where $\varepsilon$ is the clipping parameter, the likelihood ratio is defined as $\gamma_{i,k}(\theta)=\frac{\pi_{\theta}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}{\pi_{\theta_{\mathrm{old}}}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}$. Although feedback tokens $\bm{o}$ are masked out during loss computation, they still influence the context of subsequent token predictions.
4 Lazy Likelihood Displacement in Tool-integrated GRPO
Recent work (deng2025effect) introduces Lazy Likelihood Displacement (LLD) in GRPO for text-based, non-tool settings, showing that the likelihood of correct responses often decreases or improves only marginally during optimization. In this section, we extend LLD to the tool-integrated RL regime and define:
Definition 4.1 ( Tool-Lazy Likelihood Displacement)
Let $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{fin}}}$ denote the initial and finetuned policies obtained before and after optimizing a preference-learning objective $\mathcal{J}$ (e.g., Sec.˜3) over a dataset $\mathcal{D}$, with $\mathcal{J}(\theta_{\mathrm{fin}})<\mathcal{J}(\theta_{\mathrm{old}})$. Consider a tool-integrated trajectory consisting of alternating actions and feedback, $(\bm{y}_{0},\bm{o}_{0},\;\bm{y}_{1},\bm{o}_{1},\;\ldots,\;\bm{y}_{T}),$ where only the actions $\{\bm{y}_{t}\}_{t=0}^{T}$ are used in likelihood computation (feedback is masked). For each response action $\bm{y}_{t}$, define its log-likelihood change as
$\Delta_{t}(\mathbf{x},\bm{y}_{t}):=\ln\pi_{\theta_{\mathrm{fin}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t})-\ln\pi_{\theta_{\mathrm{old}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t}).$
We say that Lazy Likelihood Displacement (LLD) occurs for the action $t$ if $\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon$, where $\epsilon$ is a small or non-positive constant. And LLD occurs for the whole response if
$\sum_{t=0}^{T}\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon,$ (2)
Thus, LLD captures the failure mode where one or more response actions exhibit negligible or even negative improvement in likelihood under the optimized policy.
To better understand the origin of LLD, we highlight how negative gradients suppress the likelihood of correct actions. The following informal theorem (the formal version and proof see App.˜A) restates Theorem 4.4 from deng2025effect in the setting of tool-integrated RL:
Theorem 4.2 (Informal: Trajectory-Level LLD in Tool-Integrated GRPO)
In tool-integrated GRPO, the likelihood of an otherwise correct response can decrease when (i) incorrect responses of low likelihood and (ii) incorrect responses whose embeddings closely resemble the correct one induce large negative gradients that dominate the positive updates. These forces jointly lead to trajectory-level Lazy Likelihood Displacement.
In the setting of tool-integrated GRPO, we observed with striking frequency that the likelihood of correct responses fails to improve. This indicates an especially acute form of LLD ($\epsilon\leq 0$), causing a progressive decay in the model’s overall output likelihood. This compounding decay is a distinctive failure mode of Search-R1 style tool-integrated RL.
5 Lazy Likelihood Displacement in tool-integrated GRPO
In this section, we illustrate the prevalence of LLD in tool-integrated GRPO training and show how the LLD ($\epsilon\leq 0$) can catastrophically drive the model into training collapse.
5.1 Likelihood Dynamic
To illustrate the prevalence and progression of LLD during training, we visualize the likelihood-displacement trajectory in tool-integrated RL. As shown in Fig.˜2, the evolution unfolds in three characteristic phases. Phase I (early stagnation): the likelihood of correct responses remains nearly unchanged even as the reward increases, showing the initial emergence of LLD.
The later two phases correspond to the $\epsilon\leq 0$ regime, where the likelihood of correct responses strictly decays rather than improves. Phase II (steady-decay, steps 60–120): the likelihood exhibits a slow but consistent downward drift, while the reward grows only gradually and the gradient norm stays stable—indicating persistent LLD. Phase III (acceleration, after step 120): the likelihood begins to collapse sharply, coinciding with a rapid surge in gradient magnitude (marked by the red star), which triggers gradient explosion and ultimately leads to training collapse.
The zoomed-in view on the right highlights this transition: likelihood keeps decreasing while gradients accelerate upward, and the reward curve begins to fall. Although the final collapse is caused by exploding gradients, the likelihood decay is present throughout training, introducing an accumulating instability that becomes dramatically amplified during the acceleration phase (see Appendix Sec.˜B.1 for more discussion).
5.2 Lazy Likelihood Displacement Death Spiral
In Fig.˜2, we observe an accelerated decline in response likelihood as training progresses. In this section, we first simplify notation by writing the feedback-masked trajectory as $\hat{\bm{y}}=(\bm{y}_{0},\bm{y}_{1},\ldots,\bm{y}_{T})$. We then formally characterize this accelerated likelihood decay as a Lazy Likelihood-Displacement (LLD) death spiral.
Definition 5.1 (LLD Death Spiral)
Consider a tuple $(\bm{x},\hat{\bm{y}}^{+})$ and a policy update from $\pi_{\theta_{t}}$ to $\pi_{\theta_{t+1}}$. We say the system enters an LLD Death Spiral if the policy evolution exhibits the following self-reinforcing progression:
$\mathrm{LLD}_{t}\;\Longrightarrow\;C^{\mathrm{low}}_{t}\;\Longrightarrow\;\mathrm{LLD}_{t+1},\quad\epsilon_{t+1}<\epsilon_{t}\leq 0$
where $\mathrm{LLD}_{t}$ denotes lazy likelihood displacement at time $t$ with decreasing likelihood ($\epsilon_{t}\leq 0$), and $C^{\mathrm{low}}_{t}$ denotes low-confidence trajectories whose likelihood decreases under $\pi_{\theta_{t}}$. The transition $\mathrm{LLD}_{t}\Longrightarrow\ C^{\mathrm{low}}_{t}$ reflects that degraded likelihood leads to increasingly diffuse predictions, while the transition $C^{\mathrm{low}}_{t}\Longrightarrow\ \mathrm{LLD}_{t+1}$ occurs because low-confidence incorrect responses often contain actions similar to those in correct responses, inducing a larger value in Eq.˜9 and thus a stronger negative-gradient influence. When these effects compound across iterations, the likelihood decay accelerates, creating a self-perpetuating collapse that we refer to as the LLD Death Spiral.
Then we demonstrate the LLD Death Spiral through accelerated entropy explosion and accelerated per-sample LLD.
Figure 3: Effect of likelihood displacement across different training iterations for the Qwen2.5-3B-Instruct model. Results are computed on the first 50 samples of the training set, discarding cases where all responses are uniformly correct or uniformly incorrect. Bars below zero (orange) indicate samples whose correct responses’ likelihood decreases after training.
Accelerated Per-Sample LLD. We conduct a controlled experiment on the searching-based question answering NQ (kwiatkowski2019natural) dataset to examine how the negative gradient of GRPO affects the log-likelihood of correct responses. Using Qwen2.5-3B-Ins (yang2024qwen2), we generate 8 rollouts per question and retain only those examples that contain a mixture of correct and incorrect responses, discarding cases where all responses are uniformly correct or incorrect. To isolate the per-sample learning dynamics, we reinitialize the model parameters $\theta$ for each individual question, apply a single GRPO update to obtain $\theta^{\prime}$, and then measure the average log-likelihood change of the correct responses:
$\Delta(\bm{x}):=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}\left[\ln\pi_{\theta^{\prime}}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})-\ln\pi_{\theta}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})\right],$ (3)
for a question $\bm{x}$ with $N^{+}$ correct responses. We present the results in Fig.˜3. In the early stage of training (iteration 50), only mild LLD is observed: most questions show negligible reductions in the likelihood of correct responses, likely because correct and incorrect trajectories remain structurally dissimilar. However, as training progresses, the effect becomes more pronounced. Increasing structural similarity between incorrect and correct responses leads to broader likelihood degradation, as shown by the rising orange curve. As training enters the acceleration phase (iterations 120–140), the trend intensifies into an LD death spiral, with over half of the samples at iteration 140 showing substantial drops in the likelihood of correct responses. This collapse is driven by the extremely low likelihood of incorrect responses, which amplifies negative-gradient contributions and destabilizes learning.
Accelerated Entropy Explosion. Fig.˜4 visualizes the evolution of entropy, response length, and valid-search ratio during training for Qwen2.5-3B-Ins (a) and Qwen2.5-3B-Base (b). Across both models, mean token entropy initially increases slowly, mirroring the early slow-LLD regime in Fig.˜2, but later transitions into a phase of sharply accelerated growth that aligns with the acceleration phase of the LLD trajectory. This steepening entropy slope is a clear demonstration of the death spiral: as low-confidence responses accumulate, the model assigns increasingly diffuse, low-probability distributions to its tokens, which further strengthens LD and drives the system toward instability. Crucially, this entropy acceleration occurs while response length and valid-search counts remain nearly unchanged, confirming that the rising entropy and the corresponding likelihood decay is not caused by trajectory length or tool-query frequency, but is instead a direct manifestation of the LD effect itself.
Figure 4: We illustrate how entropy, response length, and valid-search ratio evolve during training. For both Qwen2.5-3B-Instruct (a) and Qwen2.5-3B-Base (b), entropy exhibits a accerlerrated upward trend prior to collapse indicating a strong LD issue. Meanwhile, the response length and valid-search times remain stable in the early stages but later begin to fluctuate markedly and eventually drop sharply.
5.3 Watch for correct actions embedded within incorrect responses.
Observing tool-integrated GRPO is substantially more susceptible to LLD than non-tool settings (deng2025effect). This section examines LLD at the granularity of response actions, where each action is defined by the segments between consecutive tool calls. Our analysis on Qwen2.5-3B-Ins (yang2024qwen2) reveals a striking pattern unique to tool-integrated GRPO: correct actions frequently appear within otherwise incorrect responses, and this intermixed structure disrupts the likelihood estimation process, amplifying LLD and destabilizing training.
Frequent correct actions in Incorrect responses. In our experiments, we found that the model typically uses its first response action to generate a search query, and the correctness of this initial action increases steadily as training progresses. To be specific, we measure whether the retrieved documents produced by an incorrect response match those retrieved by a correct response; if so, we treat the first action as correct. The accuracy of this initial action across different training stages is shown in Fig. 5. As indicated by the dashed green line, accuracy is low in the early stages, reflecting the model’s initial acquisition of search capability, but rises markedly over time. By step 140, roughly 60% of incorrect responses begin with a correct search query, highlighting the high structural similarity shared by first-action actions.
Figure 5: Evolution of token log-likelihood (measured before vs. after feedback; left axis) and the observation-match ratio for wrong answers (right axis). With training, both likelihoods drop while the overlap between tool observations in incorrect and correct trajectories increases, suggesting many incorrect responses begin with a correct search, which skews likelihood estimates and contributes to LLD.
Faster likelihood decay of the first-action response. As illustrated in Fig. 5, we further observe that the likelihood of the first action (light blue) declines substantially more than that of the second action (blue). Early in training, the first-action likelihood is higher, largely because the second action must condition on out-of-distribution feedback. However, as training progresses, both actions exhibit likelihood decay, with the first action eventually dropping below the second around step 110, where the first-action correctness rate is roughly 50%, indicating strong similarity between correct and incorrect trajectories. After step 120, this decay accelerates sharply, marking the onset of the LLD death spiral, where low-likelihood incorrect partial signals further amplify degradation.
Nonsensical first-action. As LLD intensifies and the likelihood of the first action collapses, the model begins producing nonsensical outputs. As illustrated in Fig.˜15 in the appendix, it eventually generates random, meaningless tokens, an effect driven by the severely reduced response likelihood, which causes sampling to select arbitrary words. Although full collapse may not yet have occurred at this stage, the model is already effectively unusable, as it can no longer produce meaningful responses.
As a result, it becomes essential to reduce the unintended penalization of correct actions within incorrect responses. A more detailed discussion is provided in Sec.˜7.
5.4 LLD regularization
To address Lazy Likelihood Displacement (LLD), we introduce a family of likelihood-preserving regularizers that prevent the model from unintentionally reducing the likelihood of responses during GRPO training. Specifically, for the given preserving responses $\bm{y}_{i}\in\mathcal{Y}_{\rm pre}$, we compare its token-level likelihood under the old policy (previous step) and the fine-tuned policy. Our base regularizer penalizes only the tokens whose likelihood decreases after an update.
LLD: Token-Level Likelihood Preservation. LLD applies a token-level penalty: within each response $\mathbf{y}_{i}$, only likelihood-reducing tokens contribute to the loss:
$L_{\rm LLD}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (4)
This token-level selectivity ensures that the model is penalized only for genuinely harmful likelihood reductions, without interfering with updates that improve the response overall. However, some responses may still be globally improving even if a few individual tokens decrease; penalizing such cases may introduce overly strong constraints.
LLDS: Response-Level Gating. To avoid unnecessary penalties on globally improving responses, LLDS introduces a response-level gating mechanism: the penalty activates only when the total likelihood of the response decreases. The LLDS loss is:
$L_{\rm LLDS}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}$ $\underbrace{\mathbbm{1}\!\left[\sum_{y_{i}\in\mathbf{y}_{i}}\left(\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)>0\right]}_{\text{Activated only when sum > 0}}\cdot$
$\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (5)
This structure preserves normal GRPO learning while directly suppressing responses that suffer from LLD.
LLDS-MA: Masking Answer Tokens. To further encourage multi-step reasoning and tool usage, we mask out final answer tokens from the regularization term. LLDS-MA penalizes likelihood reductions only on reasoning and tool-interaction tokens, excluding the answer span $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}$}:
$\underbrace{\sum_{y_{i}\in\mathbf{y}_{i}/\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}}}\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Mask Answer Likelihood-reducing tokens}}$ (6)
Finally, We integrate the regularization into the GRPO objective as
$L_{\rm total}=L_{\rm GRPO}+\lambda L_{\rm{LLDS(-MA)}}$ (7)
where $\lambda$ is the regularization weight. We use LLDS as the default variant and switch to LLDS-MA when stronger encouragement of tool usage is desired. For the preserving set $\mathbf{y}]_{\rm pre}$, we include all responses with non-negative advantages $\hat{A}\geq 0$, ensuring that correct responses ($\hat{A}>0$) and untrained responses ($\hat{A}=0$) do not suffer likelihood reduction. The effect of $\lambda$ is examined empirically in Fig.˜8.
6 Experiments and Analysis
We evaluate the empirical effectiveness of $L_{\rm LLDS(-MA)}$ through comprehensive experiments.
Experimental settings. For training, we follow the setup in jin2025search and conduct experiments using two model families: Qwen-2.5-3B and Qwen-2.5-7B, each in both Base and Instruct variants (yang2024qwen2). We consider two training configurations. (1) NQ-Only (single-hop): the model is trained solely on the single-hop Natural Questions (NQ) dataset (kwiatkowski2019natural). (2) NQ+Hotpot (single-hop+multi-hop): the model is trained on a merged corpus combining NQ with HotpotQA (yang2018hotpotqa), providing broader coverage of both open-domain and multi-hop reasoning. For the retrieval dataset, we use the 2018 Wikipedia dump (karpukhin2020dense) as the knowledge corpus and employ E5 (wang2022text) as the dense retriever. To ensure fair comparison across retrieval-augmented baselines, we fix the number of retrieved passages to three, following the configuration in jin2025search. Unless otherwise specified, we use the same optimization hyperparameters as Search-R1 (jin2025search), with the only modification being a reduced maximum turn limit: from 4 in Search-R1 to 2 for NQ-only training and 3 for the NQ+Hotpot setting, improving training efficiency. Unless stated otherwise, the regularization weight is fixed at $\lambda=0.1$.
Evaluation settings. The Evaluation is performed on the validation or test sets of seven datasets, categorized as follows: (1) General Question Answering: NQ (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa), and PopQA (mallen2022not). (2) Multi-Hop Question Answering: HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (ho2020constructing), Musique (trivedi2022musique), and Bamboogle (press2022measuring). We adopt exact match (EM) as the primary evaluation metric, consistent with the protocol in jin2025search. Such diversity ensures a thorough examination of search engine-integrated LLM performance across a range of realistic reasoning scenarios.
6.1 Experimental Results
Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
Direct Inference 0.106 0.288 0.108 0.167 0.149 0.244 0.020 0.024 0.109 0.134
CoT 0.023 0.032 0.005 0.020 0.021 0.021 0.002 0.000 0.011 0.015
IRCoT 0.111 0.312 0.200 0.208 0.164 0.171 0.067 0.240 0.161 0.181
Search-o1 0.238 0.472 0.262 0.324 0.221 0.218 0.054 0.320 0.203 0.255
RAG 0.348 0.544 0.387 0.426 0.255 0.226 0.047 0.080 0.152 0.270
SFT 0.249 0.292 0.104 0.215 0.186 0.248 0.044 0.112 0.147 0.176
R1-base 0.226 0.455 0.173 0.285 0.201 0.268 0.055 0.224 0.187 0.229
R1-instruct 0.210 0.449 0.171 0.277 0.208 0.275 0.060 0.192 0.184 0.224
Rejection Sampling 0.294 0.488 0.332 0.371 0.240 0.233 0.059 0.210 0.186 0.265
Search-R1-PPO-Base 0.406 0.587 0.435 0.476 0.284 0.273 0.049 0.088 0.174 0.303
Search-R1-PPO-Ins 0.341 0.545 0.378 0.421 0.324 0.319 0.103 0.264 0.253 0.325
Qwen2.5-3b-Base
NQ-Only
Search-R1-GRPO 0.440 0.582 0.413 0.478 0.265 0.244 0.061 0.113 0.171 0.303
+LLD 0.462 0.609 0.464 0.512 0.286 0.248 0.063 0.113 0.178 0.321 (+5.9%)
+LLDS 0.478 0.605 0.449 0.511 0.288 0.250 0.062 0.129 0.182 0.323 (+6.6%)
NQ+hotpot
Search-R1-GRPO 0.421 0.583 0.413 0.472 0.297 0.274 0.066 0.128 0.191 0.312
+LLDS 0.479 0.634 0.462 0.525 0.350 0.308 0.086 0.201 0.236 0.360 (+15.4%)
+LLDS-MA 0.486 0.632 0.462 0.527 0.444 0.430 0.183 0.371 0.357 0.430 (+37.8%)
Qwen2.5-3b-Ins
Search-R1-GRPO 0.397 0.565 0.391 0.451 0.331 0.310 0.124 0.232 0.249 0.336
NQ-Only
+LLDS 0.490 0.610 0.450 0.517 0.294 0.193 0.074 0.121 0.171 0.319
+LLDS-MA 0.464 0.616 0.458 0.513 0.363 0.320 0.131 0.315 0.282 0.381 (+13.4%)
NQ+hotpot
+LLDS 0.462 0.622 0.460 0.515 0.432 0.383 0.180 0.395 0.347 0.419 (+24.7%)
Table 1: Results on General QA and Multi-Hop QA datasets for Qwen2.5-3b-Base/Instruct. denotes results from (jin2025search) using NQ+hotpot training. denotes open-domain datasets and denotes retrieval-augmented datasets.
We evaluate our proposed LLD-mitigation strategy on seven open-domain and multi-hop QA benchmarks using two model families: Qwen2.5-3B (Base/Instruct) and Qwen2.5-7B (Base/Instruct). Tables 1 and 2 summarize the EM performance across all settings. Since vanilla GRPO training often collapses and drives the reward to zero, we use the results from Search-R1 jin2025search, corresponding to the best checkpoint before collapse, as the GRPO baseline.
Results on Qwen2.5-3B. As shown in Tab.˜1, GRPO attains an average score of 0.303 in the NQ-only setting. Incorporating LLD raises performance to 0.321 (+5.9%), and LLDS provides a further improvement to 0.323 (+6.6%). In the more challenging NQ+Hotpot setting, LLDS increases the vanilla GRPO score from 0.312 to 0.360, a 15.4% relative gain. Notably, since LLDS tends to limit the base model to invoking search only once, we apply the multi-aspect variant LLDS-MA, which delivers the strongest performance with an average score of 0.430, representing a substantial +37.8% improvement ratio over vanilla GRPO. A similar trend is observed for Qwen2.5-3B-Instruct: LLDS-MA improves performance to 0.381 under NQ-only, and LLDS achieves 0.419 in the NQ+Hotpot setting (+24.7%).
Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
Qwen2.5-7b-Base/Instruct
Direct Inference 0.134 0.408 0.140 0.227 0.183 0.250 0.031 0.120 0.146 0.181
CoT 0.048 0.185 0.054 0.096 0.092 0.111 0.022 0.232 0.114 0.106
IRCoT 0.224 0.478 0.301 0.334 0.133 0.149 0.072 0.224 0.145 0.239
Search-o1 0.151 0.443 0.131 0.242 0.187 0.176 0.062 0.296 0.180 0.206
RAG 0.349 0.585 0.392 0.442 0.299 0.235 0.058 0.208 0.200 0.304
SFT 0.318 0.354 0.121 0.264 0.217 0.259 0.066 0.112 0.164 0.207
R1-base 0.297 0.539 0.199 0.345 0.242 0.273 0.083 0.203 0.200 0.262
R1-instruct 0.270 0.537 0.199 0.335 0.237 0.292 0.072 0.293 0.224 0.271
Rejection Sampling 0.360 0.592 0.380 0.444 0.331 0.296 0.123 0.355 0.276 0.348
Search-R1-PPO-Base 0.480 0.638 0.457 0.525 0.433 0.382 0.196 0.432 0.361 0.431
Search-R1-PPO-Ins 0.393 0.610 0.397 0.467 0.370 0.414 0.146 0.368 0.325 0.385
Qwen2.5-7b-Base
Search-R1 (GRPO) 0.395 0.560 0.388 0.448 0.326 0.297 0.125 0.360 0.277 0.350
NQ-Only
+LLDS 0.514 0.666 0.470 0.550 0.403 0.368 0.151 0.347 0.317 0.417 (+19.1%)
NQ+hotpot
+LLDS 0.504 0.666 0.471 0.547 0.465 0.475 0.209 0.444 0.398 0.462 (+32.0%)
Qwen2.5-7b-Ins
Search-R1 (GRPO) 0.429 0.623 0.427 0.493 0.386 0.346 0.162 0.400 0.324 0.396
NQ-Only
+LLDS 0.517 0.645 0.460 0.541 0.401 0.298 0.158 0.411 0.317 0.413 (+4.3%)
NQ+hotpot
+LLDS 0.498 0.662 0.464 0.541 0.473 0.443 0.228 0.516 0.428 0.469 (+18.4%)
Table 2: Performance comparison on General QA and Multi-Hop QA datasets for Qwen2.5-7b-Base/Instruct.
Results on Qwen2.5-7B As shown in Tab.˜2, for Qwen2.5-7B-Base, applying LLDS yields a substantial improvement, reaching an average score of 0.417 when trained on NQ only, representing a 19.1% relative gain over the GRPO baseline trained on NQ+Hotpot. When LLDS is applied under the NQ+Hotpot training setting, performance increases from 0.350 to 0.462, corresponding to an even larger 32.0% improvement. LLDS also achieves the strongest results across nearly all individual datasets, indicating enhanced evidence retrieval and multi-step reasoning capability. For Qwen2.5-7B-Instruct, we observe a similar pattern. Under the NQ+Hotpot training setting, LLDS raises the average score from 0.396 to 0.469 (improved 18.4%). Notably, LLDS achieves the best performance on every Multi-Hop QA benchmark, including 0.473 on 2Wiki, 0.443 on Musique, and 0.516 on Bamboogle.
These results highlight LLDS(-MA) as an effective method for stabilizing GRPO training and substantially enhancing multi-turn reasoning.
6.2 Ablation Study and Analysis
Impact of response-level gating. Applying response-level gating reduces the strength of the regularization applied during training, leading to measurable improvements on multi-hop QA tasks. As shown in Tab.˜1, although the average performance increases modestly by 0.2%, the method yields a notable 1.6% gain on the Bamboogle dataset, highlighting its effectiveness in enhancing multi-hop reasoning.
NQ vs. NQ+Hotpot. Models trained only on NQ achieve strong performance on open-domain, single-hop QA but show limited ability to generalize to multi-hop settings. Moving to NQ+Hotpot consistently improves multi-step reasoning: for example, on Qwen2.5-3B-Base, the GRPO baseline increases from 0.303 (NQ-only) to 0.312 (NQ+Hotpot), and LLDS increases from 0.323 to 0.360 under the same shift. A similar trend appears for Qwen2.5-7B-Base, where LLDS rises from 0.417 (NQ-only) to 0.462 (NQ+Hotpot). These gains show that adding multi-hop HotpotQA, whose questions require multi-turn reasoning to answer correctly and receive reward, encourages the model to retrieve, integrate, and reason over multiple pieces of evidence more effectively than training on NQ alone.
Impact of masking answer (MA). We apply masking answer (MA) specifically to model variants where vanilla GRPO and LLDS degenerate into issuing only a single search call, limiting their ability to collect sufficient evidence for multi-hop reasoning. MA disables LLDS regularization on the final answer tokens, encouraging the model to perform additional search steps or intermediate reasoning. As shown in Tab.˜1, adding MA on top of LLDS yields substantial improvements for these one-search–limited models. LLDS-MA increases the Qwen2.5-3B-Base average score in the NQ+Hotpot setting from 0.360 (LLDS) to 0.430, and similarly improves the NQ-only Instruct variant from 0.319 to 0.381. These gains demonstrate that MA effectively encourages deeper, multi-step reasoning in scenarios where the underlying GRPO policy underutilizes search actions. More detailed analysis see Sec.˜B.3.
6.3 Effect of LLDS on Training Stability across Models
To verify the universality of the training collapse issue in GRPO and the robustness of our proposed solution, we extended our experiments across different model scales and alignment stages. Fig.˜6 illustrates the training reward curves for Qwen-2.5 3B and 7B models, covering both Base and Instruct versions. As observed in the blue curves, the baseline GRPO consistently suffers from catastrophic collapse within the first 300 steps, regardless of the model size (3B vs. 7B) or type (Base vs. Instruct). This confirms that the instability is an inherent characteristic of the algorithm rather than an artifact of a specific model configuration. In contrast, the integration of LLDS (red curves) effectively mitigates this issue. In all four scenarios, the models trained with LLDS maintain a steady upward trajectory in rewards, successfully bypassing the collapse points that affect the baseline. This demonstrates that LLDS serves as a generalized and effective regularizer for stabilizing GRPO training.
(a) Qwen-2.5-3B-Base
(b) Qwen-2.5-3B-Instruct
(c) Qwen-2.5-7B-Base
(d) Qwen-2.5-7B-Instruct
Figure 6: Comparisons of training reward curves between baseline GRPO (blue) and GRPO + LLDS (red). The baseline consistently collapses (reward drops to zero) across all model sizes, while LLDS stabilizes training and sustains high rewards.
7 More Discussion and Guideline
Based on our analysis of Lazy Likelihood Displacement (LLD) and the emergence of the LLD Death Spiral, we consolidate several practical guidelines for stabilizing tool-integrated GRPO. Each recommendation follows directly from the core failure mechanisms identified in our study.
Understand why tool-integrated GRPO is uniquely vulnerable to LLD. Unlike free-form RL on text, tool-augmented agents introduce structural conditions that magnify likelihood drift. First, tool calls inject inherently out-of-distribution tokens—such as search results, API outputs, or error messages—that differ sharply from pretrained language distributions. These OOD segments raise token-level uncertainty and make GRPO’s relative updates more volatile, accelerating the onset of likelihood displacement. Second, tool-based reasoning unfolds across multiple stages, with early stages (e.g., query formulation) stabilizing more quickly than later ones (e.g., tool-result interpretation). Because GRPO applies a single scalar reward to all tokens, early-stage tokens that share nearly identical prefixes across correct and incorrect trajectories receive conflicting gradient signals. This reward–token misalignment disproportionately harms stable prefixes and amplifies LLD. Recognizing these structural challenges is essential for diagnosing and mitigating instability in tool-integrated RL systems.
Closely monitor likelihood dynamics—reward alone is insufficient. A central insight of our analysis is that likelihood degradation begins long before any visible drop in reward. Across models, reward continues to rise throughout both the early and steady-decay phases, even as the likelihood of correct responses is already drifting downward. Because GRPO updates are governed by likelihood ratios, such early degradation silently amplifies gradients and initiates instability. Monitoring only the reward therefore obscures the onset of LLD. In practice, reliable early-warning signals come from tracking action-level and trajectory-level log-likelihood, inspecting entropy trajectories, which surge sharply once likelihood begins to collapse, and watching for sudden spikes in likelihood ratios that indicate gradient amplification. Visualization of likelihood or entropy curves provides the earliest and most trustworthy indication that training is approaching the unstable regime.
8 Conclusion
In this work, we firstly identified Lazy Likelihood Displacement as the central cause of instability in GRPO-based, Search-R1–style tool-integrated RL. We showed that LLD emerges early, occurs frequently, and reliably drives training into a self-reinforcing LLD Death Spiral characterized by monotonic likelihood decay, entropy spikes, gradient inflation, and eventual collapse. Our analysis provides the first systematic evidence that GRPO is uniquely vulnerable to likelihood drift in tool-integrated settings. To counteract this failure mode, we proposed $\rm LLDS$, a simple and effective likelihood-preserving regularizer that activates only when likelihood decreases and targets only the offending tokens. Empirically, $\rm LLDS$ stabilizes optimization and yields consistent gains across seven QA benchmarks, with improvements of up to +37.8% on Qwen2.5-3B and +32.0% on Qwen2.5-7B. These findings highlight likelihood control as a key mechanism for reliable RL-based tool use and offer a promising foundation for more robust and scalable training methods in future agentic LLMs.
Acknowledgments: This work was partially funded by the NSERC Discovery Grant RGPIN-2021-03677, Alliance Grant ALLRP 581098-22, the Natural Science and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs program, the Canada Research Chair program, an IITP grant funded by MSIT, and the Digital Research Alliance of Canada.
Appendix A Theorem and Proof
In this section, we first give the formal version of  theorem˜4.2:
Theorem A.1 (Trajectory-Level LLD in Tool-Integrated GRPO)
Consider a tool-integrated trajectory with only the response actions $\{\bm{y}_{t}\}_{t=0}^{T}$ contribute to the likelihood, and tool-feedback tokens $\bm{o}_{t}$ are masked during likelihood computation but remain in context. Let $\pi_{\theta(s)}$ denote the evolving policy at training time $s$.
action-level likelihood change.
[…]
Trajectory-level likelihood change.
Summing over all actions yields $\frac{d}{ds}\ln\pi_{\theta(s)}\!\left(\bm{y}_{0:T}\mid\bm{x}\right)=\sum_{t=0}^{T}\sum_{k=1}^{|\hat{\bm{y}}_{t}^{+}|}\frac{d}{ds}\ln\pi_{\theta(s)}(\hat{\bm{y}}_{t,k}^{+}|\cdot),$ which becomes lazy or negative whenever $\sum_{t=0}^{T}\sum_{k=1}^{|\hat{\bm{y}}_{t}^{+}|}\mathcal{G}_{t,k}(s)\quad\text{is large}.$
As the theorem indicates, two core factors inflate this negative-gradient effect:
  1. 1.
    Low likelihood of incorrect responses: Negative responses that the model assigns low probability to yield larger prediction-error weights $\alpha^{-}_{t,k;\,t^{\prime},k^{\prime}}$ , magnifying their influence. In such cases, the model interprets these low-likelihood errors as severe mistakes, causing their gradients to receive disproportionately large scaling.
  2. 2.
    Embedding similarity: When incorrect responses are similar to correct ones, their prefix representations exhibit large inner products, amplifying the negative contribution. This high representational overlap means the model struggles to disentangle correct from incorrect continuations, causing negative examples to push gradients in harmful directions and make the model unconfident.
A.1 Proof of  theorem˜A.1
Setup and masking.
Fix a query $\bm{x}$ and a correct response index $i$, and consider the feedback-masked trajectory
[…]
where only the action tokens in $\{\bm{y}^{+}_{i,t}\}_{t=0}^{T_{i}}$ contribute to the loss. Tool feedback $\bm{o}$ is excluded from the GRPO objective but remains in the conditioning context. We study the log-likelihood change of each action $\bm{y}^{+}_{i,t}$ under the evolving policy $\pi_{\theta(s)}$:
[…]
and then aggregate over $t$ to obtain the trajectory-level result.
Reduction to standard GRPO at the action level.
[…]
Since feedback tokens are only masked in the loss, but still appear in the context, the GRPO objective for tool-integrated training (with feedback masking) has exactly the same functional form as the standard GRPO objective applied to the sequence of action tokens. Thus, each pair
[…]
Theorem A.2 (Action-level)
For any $\bm{x}$, any time $s\geq 0$, and any correct response $\bm{y}_{i}^{+}$, the likelihood change of its $t$-th action,
[…]
becomes lazier (smaller in magnitude, and potentially negative) as
$\mathcal{G}_{i,t}(s)$ $=\underbrace{p^{-}\sum_{k=0}^{|\bm{y}^{+}_{i,t}|}\sum_{j=1}^{N^{-}}\sum_{t^{\prime}=0}^{T_{j}}\sum_{k^{\prime}=1}^{|\bm{y}_{j,t^{\prime}}^{-}|}\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}\,\bigl\langle\mathbf{h}_{\bm{x},\bm{y}^{+}_{i,<t},\bm{o}^{+}_{i,<t},\bm{y}_{i,t,<k}^{+}},\mathbf{h}_{\bm{x},\bm{y}^{-}_{j,<t^{\prime}},\bm{o}^{-}_{j,<t^{\prime}},\bm{y}_{j,t^{\prime},<k^{\prime}}^{-}}\bigr\rangle}_{\text{impact of negative gradients}}$
$\quad-p^{+}\sum_{k=0}^{|\bm{y}^{+}_{i,t}|}\sum_{i^{\prime}=1}^{N^{+}}\sum_{t^{\prime\prime}=0}^{T_{i^{\prime}}}\sum_{k^{\prime\prime}=1}^{|\bm{y}_{i^{\prime},t^{\prime\prime}}^{+}|}\alpha^{+}_{(i,t,k),(i^{\prime},t^{\prime\prime},k^{\prime\prime})}\,\bigl\langle\mathbf{h}_{\bm{x},\bm{y}^{+}_{i,<t},\bm{o}^{+}_{i,<t},\bm{y}_{i,t,<k}^{+}},\mathbf{h}_{\bm{x},\bm{y}^{+}_{i^{\prime},<t^{\prime\prime}},\bm{o}^{+}_{i^{\prime},<t^{\prime\prime}},\bm{y}_{i^{\prime},t^{\prime\prime},<k^{\prime\prime}}^{+}}\bigr\rangle.$ (9)
increases, where
$\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}=\left\langle\mathbf{e}_{y^{+}_{i,t,k}}-\pi_{\theta(s)}(\cdot\mid x,y^{+}_{i,<t},o^{+}_{i,<t},y^{+}_{i,t,<k}),\mathbf{e}_{y^{-}_{j,t^{\prime},k^{\prime}}}-\pi_{\theta(s)}(\cdot\mid x,y^{-}_{j,<t^{\prime}},o^{-}_{j,<t^{\prime}},y^{-}_{j,t^{\prime},<k^{\prime}})\right\rangle$ and
[…]
Trajectory level as a sum over actions.
The tool-integrated trajectory likelihood factorizes over actions:
$\pi_{\theta(s)}(\bm{y}_{0:T}\mid\bm{x})=\prod_{t=0}^{T}\pi_{\theta(s)}\!\left(\bm{y}_{t}\mid\bm{x},\bm{y}_{<t},\bm{o}_{<t}\right).$
Taking logarithms and differentiating with respect to $s$ gives
$\frac{d}{ds}\ln\pi_{\theta(s)}(\bm{y}_{0:T}\mid\bm{x})$ $=\sum_{t=0}^{T}\frac{d}{ds}\ln\pi_{\theta(s)}\!\left(\bm{y}_{t}\mid\bm{x},\bm{y}_{<t},\bm{o}_{<t}\right)$
[…]
where $\cdot$” again denotes the masked context including the feedback tokens. By the action-level result, each term in the sum becomes lazy or negative as $\mathcal{G}_{t,k}(s)$ grows, and hence the full trajectory-level derivative becomes lazy or negative whenever
$\sum_{t=0}^{T}\sum_{k=1}^{|\hat{\bm{y}}_{t}^{+}|}\mathcal{G}_{t,k}(s)$
is large. This establishes the trajectory-level statement in Theorem˜A.1.
Appendix B Additional Analysis
B.1 Training Instability Prior to Collapse
Figure 7: Demonstration of unstability prior collapse.
While the final collapse is indeed triggered by large gradients from negative responses, substantial instability and a noticeable performance decline had already emerged earlier in training. As shown in Fig.˜7, the brown vertical line marks the first sharp drop in model reward (green curve), during which the likelihood of correct responses also decreases significantly, even though the gradient norm remains relatively small (around 2). As training progresses, the LD issue intensifies and the low confidence responses lead to gradient norm steadily grows, ultimately causing the model to collapse. These observations suggest that instability is present well before the collapse point and therefore warrants close attention.
B.2 Impact of Regularization Strength
To examine the effect of regularization strength, we ablate $\lambda\in{0,0.01,0.1}$ and plot the corresponding training reward dynamics in Fig.˜8. As shown, without regularization ($\lambda=0$), the model collapses around step 200. A small regularization value ($\lambda=0.01$, orange curve) delays but does not prevent collapse, which occurs around step 220. In contrast, a stronger regularization ($\lambda=0.1$) stabilizes training entirely, allowing the model to continue smoothly without collapse. We truncate the plot at step 250 for clarity, although training can proceed well beyond this point.
Figure 8: Impact of regularization strength on Qwen2.5-7B-base.
B.3 Unlocking Multi-step Reasoning via Answer Masking
To further explore the potential of LLDS in promoting complex reasoning behaviors, we conducted an ablation study on the Qwen-2.5-3B-Base model, focusing on the regularization scope. Specifically, we mask out the final answer tokens from the calculation of the $L_{\text{LLDS}}$ term (denoted as "Mask Answer").
It is important to note that Qwen-2.5-3B-Base inherently lacks the capability for multi-turn tool calling. As illustrated in Fig.˜9, this limitation is evident in the standard training setups: the baseline GRPO (blue curve) rapidly suffers from model collapse, with the number of valid searches dropping to zero. While the standard GRPO + LLDS (red curve) successfully stabilizes the training and prevents collapse, it fails to elicit any multi-step behavior, with the search frequency remaining stagnant at exactly 1.0.
However, the "GRPO + LLDS-MA" variant (green curve) demonstrates a distinct emergent behavior. By removing the regularization constraint on the final answer tokens, we observe a significant increase in the number of valid searches per question, rising well above 2.0. This indicates that relaxing the penalty on the answer tokens effectively unlocks the model’s latent potential to engage in multi-step reasoning and utilize external tools—capabilities that standard methods failed to activate in the base model.
Figure 9: Evolution of the number of valid searches per question on Qwen-2.5-3B-Base. The base model inherently lacks multi-turn tool calling capabilities. While the baseline (blue) collapses and standard LLDS (red) stabilizes at a single search step, LLDS-MA (green) successfully unlocks the model’s potential for multi-step reasoning, driving the number of valid searches significantly above 1.0.
B.4 Qualitative Examples of Tool-Integrated Reasoning
To further illustrate the behavioral characteristics of our trained models, we present qualitative reasoning traces generated by the Qwen2.5-7B family. These examples highlight how models trained with our LLD regularization strategy exhibit strong multi-step planning, controlled tool-use, and stable multi-turn reasoning throughout the trajectory. Unlike standard GRPO models—which often suffer from premature collapse, our method enables the model to maintain coherent reasoning structure across search, verification, and final answer generation. As shown in Figures 10 and 11, the model not only performs step-by-step decomposition and self-verification but also executes consecutive search calls when needed, integrates retrieved evidence effectively, and produces a correct, concise final answer. These qualitative behaviors align with our quantitative findings: stabilizing likelihood dynamics prevents the LLD Death Spiral, allowing the RL policy to leverage deeper tool-integrated reasoning without sacrificing robustness.
Appendix C Case Studies of Likelihood Displacement
To better understand how Lazy Likelihood Displacement manifests in practice, we zoom into two representative training samples (corresponding to the most severely affected two samples in Fig. 3 (Step 140)). For each question, we compare a correct trajectory with an incorrect one generated by the SEARCH-R1 policy. These paired examples make the abstract mechanisms in our analysis concrete.
Case 1: Embedding similarity under group-relative updates. Fig. 12 and Fig. 13 show two rollouts for the question “Who won the NRL grand final in 2015?”. Both trajectories use nearly identical <think> plans, issue semantically equivalent search queries, and retrieve identical <information> snippets. The only difference lies in the final answer token: the correct trajectory outputs the full entity “North Queensland Cowboys”, whereas the incorrect one truncates the name to “North Queensland”. Because GRPO assigns a single scalar reward to the entire trajectory, these two highly similar prefixes receive sharply different updates: the incorrect rollout pushes negative gradients onto tokens that are almost identical in embedding space to those of the correct rollout. This illustrates how small semantic deviations at the end of a trajectory can, through high prefix similarity, induce strong likelihood displacement on otherwise correct actions.
Case 2: Low-likelihood, longer incorrect trajectories. Fig. 14 and Fig. 15 present rollouts for “Who is the main character in green eggs and ham?”. In the incorrect case, the model begins with a long, low-likelihood <think> segment. Because the likelihood of early tokens is extremely small, sampling drifts into semantically meaningless regions of the distribution, producing an extended sequence of nonsensical text. This also leads the model to violate the tool protocol, which subsequently triggers corrective feedback in the <information> channel. Although the model eventually issues a valid search, it still produces an incorrect answer (“The first-person narrator”). This trajectory is both longer and lower-likelihood than its correct counterpart, causing GRPO to assign it large negative prediction-error weights and accumulate many negative-gradient summands across tokens. In contrast, the correct trajectory remains short and high-confidence, performing a single search followed by the correct answer “Sam-I-Am”. Together, these examples concretely demonstrate how low-likelihood, overlong incorrect responses can dominate gradients and drive the LLD Death Spiral, even when the environment occasionally nudges the model back toward valid tool use.
Question: Who was the screenwriter for Three Times? Ground Truth: Hou Hsiao-hsien SEARCH-R1: <think>
To answer this question, I need to find out who the screenwriter for the movie Three Times is. I will conduct a search to gather this information.
</think> <search> who was the screenwriter for Three Times </search> <information>
Doc 1 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
Doc 2 (Title: Paddy Chayefsky) Paddy Chayefsky Sidney Aaron "Paddy" Chayefsky (January 29, 1923 – August 1, 1981) was an American playwright
Doc 3 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
</information>
<think>
From the search results, I found that the director of Three Times is Hou Hsiao-hsien. However, I need to find the screenwriter. Since the search results did not provide this information, I will conduct another search.
</think> <search> screenwriter of Three Times </search> <information>
Doc 1 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
Doc 2 (Title: Eric Mendelsohn) Eric Mendelsohn (born November 1, 1964) is an American film director and screenwriter. Two of his films have been screened
Doc 3 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
</information>
<think>
From the search results, I found that the director of Three Times is Hou Hsiao-hsien. However, I need to find the screenwriter. Since the search results did not provide this information, I will infer that the screenwriter is Hou Hsiao-hsien, as he is the director of the film.
</think> <answer> Hou Hsiao-hsien </answer>
Figure 10: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search, with self-verification and finally answer the question.
Question: Which was founded first University of California Santa Barbara or Hamdard University? Ground Truth: University of California, Santa Barbara SEARCH-R1: <think>
To answer this question, I need to find out which university was founded first. I’$l$ conduct a search to gather this information.
</think> <search> Which was founded first, University of California, Santa Barbara or Hamdard University </search> <information>
Doc 1(Title: "University of California, Santa Barbara") University of California, Santa Barbara The University of California, Santa Barbara
Doc 2(Title: "History of the University of California, Santa Barbara") History of the University of California, Santa Barbara The University of California, Santa Barbara (UCSB)
Doc 3(Title: "History of the University of California, Santa Barbara") the chief executive was changed to Chancellor. UCSB’s first Provost was thus Clarence L.
</information>
<think>
From the search results, I found that the University of California, Santa Barbara traces its roots back to 1891. Now, I need to search for the founding date of Hamdard University.
</think> <search> Hamdard University founding date </search> <information>
Doc 1(Title: "Hamdard University") century. The university includes eight faculties, nine research institutes, three teaching hospitals and three affiliated engineering institutes.
Doc 2(Title: "Hamdard University") Hamdard University Hamdard University (Urdu: ) () is a private research university with campuses in Karachi
Doc 3(Title: "Hamdard University") electrical engineer Atta-ur-Rehman Memon, and Zillur Rahman (who has been serving as visiting professor of medicine since 1997)
</information>
<think>
From the search results, I found that Hamdard University was founded in 1991. Now, I can compare the founding dates of the two universities. The University of California, Santa Barbara was founded in 1891, while Hamdard University was founded in 1991. Therefore, the University of California, Santa Barbara was founded first.
</think> <answer> University of California, Santa Barbara </answer>
Figure 11: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search and finally answer the question.
Question: Who won the NRL grand final in 2015? Ground Truth: North Queensland Cowboys SEARCH-R1:
<think>
To direct to the original accessible information efficiently, let me consult the query through search.
</think>
<search>
NRL grand final 2015 winner
</search>
<information>
Doc 1(Title: "2015 NRL Grand Final") 2015 NRL Grand Final The 2015 NRL Grand Final was the conclusive and premiership-deciding game of the 2015 NRL season and was played on Sunday 4 October at Sydney’s ANZ Stadium betwee
Doc 2(Title: "2015 NRL Grand Final") grand finalists, North Queensland and Brisbane, as well as minor premiers the Sydney Roosters, qualified for the second World Club Series, played in the 2016 pre-season against clubs Doc 3(Title: "2015 NRL Grand Final") was the first time an NRL Grand Final had been decided in Golden Point. At the post-match ceremony, Johnathan Thurston received the Clive Churchill medal as best on ground. This was t
</information>
<answer>
North Queensland Cowboys
</answer>
Figure 12: Case 1 (correct). SEARCH-R1 produces a compact reasoning trace and the fully correct entity name for the NRL grand final question.
Question: Who won the NRL grand final in 2015? Ground Truth: North Queensland Cowboys SEARCH-R1:
<think>
I need hunting and insight into history of NRL Grand Final 2015. Bespoken search seem to be fitting solution here which I’$l$ commence.
</think>
<search>
who won nrl grand final 2015
</search>
<information>
Doc 1(Title: "2015 NRL Grand Final") 2015 NRL Grand Final The 2015 NRL Grand Final was the conclusive and premiership-deciding game of the 2015 NRL season and was played on Sunday 4 October at Sydney’s ANZ Stadium betwee
Doc 2(Title: "2015 NRL Grand Final") grand finalists, North Queensland and Brisbane, as well as minor premiers the Sydney Roosters, qualified for the second World Club Series, played in the 2016 pre-season against clubs Doc 3(Title: "2015 NRL Grand Final") title for the St George Illawarra Dragons in 2010. The North Queensland Cowboys reached the premiership deciding game for the second time in their history since their inception into t
</information>
<answer>
North Queensland
</answer>
Figure 13: Case 1 (incorrect). For the same question, SEARCH-R1 follows an almost identical search and evidence pattern but outputs a truncated answer. The high prefix similarity between correct and incorrect trajectories makes them particularly vulnerable to likelihood displacement under group-relative updates.
Question: Who is the main character in green eggs and ham? Ground Truth: Sam-I-am SEARCH-R1:
<think>
I’m worried I do not know tale specifically - therefore to ascertain what the main character in Green Eggs and Ham would be, I shall seek this information thusly.
</think>
<search>
who is the main character in Green Eggs and Ham
</search>
<information>
Doc 1(Title: "Green Eggs and Ham") Green Eggs and Ham Green Eggs and Ham is a children’s book by Dr. Seuss, first published on August 12, 1960. As of 2016, the book has sold 8 million copies worldwide. The story has appe
Doc 2(Title: "Green Eggs and Ham (TV series)") Green Eggs and Ham (TV series) Green Eggs and Ham is an upcoming American animated television series from Warner Bros. Animation, A Very Good Production, A Stern Talking To, Doc 3(Title: "Green Eggs and Ham (TV series)") Sam-I-Am and his friend go on a cross-country trip that would test the limits of their friendship. As they learn to try new things, they find out what adventure brings. Gree
</information>
<answer>
Sam-I-Am
</answer>
Figure 14: Case 2 (correct). A stable SEARCH-R1 trajectory for the same question: the model uses a short, coherent <think> step, performs a single search, and outputs the correct answer “Sam-I-Am”.
Question: Who is the main character in green eggs and ham? Ground Truth: Sam-I-am SEARCH-R1:
<think>
In order to resolve who the main character within the text "Green Eggs and Ham", I can infer from reading this poetic curiously imaginative nature of tales associated with Dr. Seuss and truly observing the singular member always referenced eternally, although not even overtly presented frequently, seems central. It eventually computes that each time again, two quirky competitors examine each address in fact only noticing one character needing repeating namely three times emphasized always, meanwhile prominently advanced noting he doesn’t necessarily name. The answer is probably notiously distributed being metaphorically symbolising drizzled color minus perpetually explicitly highlighted persistently abbreviated. Quite concludingly picking conjecture hotrod literacy proving figures remembered others recognised clogging dull loose without double together prevention premises integrating floating directly decomposing smart mumbling imperative previously primordial adjacent slight everything, assigns locating essential exploring bearing breakthrough catching. Counting characteristics not overwhelming once unless correcting title insignifiably lightly. Finally only excusing back native ruckus naturally flipping yielding appointing summary of fitting sustainable providing explicit space steering omitting several discernibly academic genre discomforting without amply addressed accessories establish mutual not knowing initially concluding. Rightly summing up rudimentarily realizing computing Boolean cutely delayed indeed persuading predictable binding nutrients distinguishing vocabularily interrelational strategy cognizant broad range indicating integrating equally locating essence wholly sectioning open mainly umbrage confirming linguistically certainty personistically figuring hyperlinks meeting melodic benders incentives predictive spanning reunite nearest that strictly wholly consistent factual reformaturing calculating empirically traversing parameters meeting aesthetic
<information>
My previous action is invalid. If I want to search, I should put the query between <search> and </search>. If I want to give the final answer, I should put the answer between <answer> and </answer>. Let me try again.
</information>
<search>
who is the main character in green eggs and ham
</search>
<information>
Doc 1(Title: "Green Eggs and Ham") Green Eggs and Ham Green Eggs and Ham is a children’s book by Dr. Seuss, first published on August 12, 1960. As of 2016, the book has sold 8 million copies worldwide. The story has appe
Doc 2(Title: "Green Eggs and Ham (TV series)") Green Eggs and Ham (TV series) Green Eggs and Ham is an upcoming American animated television series from Warner Bros. Animation, A Very Good Production, A Stern Talking To, Doc 3(Title: "Green Eggs and Ham (TV series)") Sam-I-Am and his friend go on a cross-country trip that would test the limits of their friendship. As they learn to try new things, they find out what adventure brings. Gree
</information>
<answer>
The first-person narrator
</answer>
Figure 15: Case 2 (incorrect). SEARCH-R1 generates a long, low-likelihood and partially nonsensical <think> segment, violates the tool protocol, receives corrective feedback, and ultimately outputs an incorrect answer. The low likelihood and extended length of this trajectory make its negative gradients particularly dominant.

### v2

@@
-On GRPO Collapse in Search-R1: The Lazy Likelihood-Displacement Death Spiral
-Abstract
-Tool-integrated (TI) reinforcement learning (RL) enables large language models (LLMs) to perform multi-step reasoning by interacting with external tools such as search engines and retrievers.
-Group Relative Policy Optimization (GRPO), exemplified by the recent Search-R1 (jin2025search), offers fast convergence and a value-free formulation that makes it appealing for this setting, yet consistently suffers from training collapse.
-We identify Lazy Likelihood Displacement (LLD), a systematic reduction or stagnation in the likelihood of both correct and incorrect responses, as the core mechanism driving this failure.
+On Group Relative Policy Optimization Collapse in Agent Search: The Lazy Likelihood-Displacement
+Abstract.
+Tool-integrated (TI) reinforcement learning (RL) enables large language models (LLMs) to perform multi-step reasoning by interacting with external tools such as search engines.
+Group Relative Policy Optimization (GRPO), exemplified by the recent Search-R1 (Jin et al., 2025), offers fast convergence and a value-free formulation that makes it appealing for this setting, yet consistently suffers from training collapse.
+We identify Lazy Likelihood Displacement (LLD), a systematic reduction or stagnation in the likelihood of correct response trajectories, as the core mechanism driving this failure.
 LLD emerges early and triggers a self-reinforcing LLD Death Spiral, where declining likelihood leads to low-confidence responses, inflating gradients and ultimately causing collapse.
-We empirically characterize this process across models on a Search-R1-style, search-integrated question answering task, revealing a consistent three-phase trajectory: early stagnation, steady decay, and accelerated collapse.
-To address this, we propose a lightweight likelihood-preserving regularization $\rm LLDS$ for GRPO that activates only when a trajectory’s likelihood decreases, and regularize only the tokens responsible.
-This fine-grained structure mitigates LLD with minimal interference to optimization.
-Across seven open-domain and multi-hop QA benchmarks, our method stabilizes training, prevents gradient explosion, and yields substantial performance improvements, including +37.8% gains on Qwen2.5-3B and +32.0% gains on Qwen2.5-7B.
-Our results establish LLD as a fundamental bottleneck in GRPO-based TIRL and provide a practical path toward stable, scalable training of tool-integrated LLM.
+We empirically characterize this process across models on a search-integrated question answering task, revealing a three-phase trajectory: early stagnation, steady decay, and accelerated collapse.
+To address this, we propose a likelihood-preserving regularization $\rm LLDS$ that activates only when a response action’s likelihood decreases, and regularizes only the tokens responsible.
+This fine-grained structure mitigates LLD with minimal interference.
+Our method stabilizes training, prevents gradient explosion, and yields substantial performance improvements across seven benchmarks, including relative improvements of +45.2% on Qwen2.5-3B and +37.1% on Qwen2.5-7B over vanilla GRPO training.
+Our results establish LLD as a previously overlooked bottleneck in GRPO-based TIRL and provide a practical path toward stable, scalable training of tool-integrated RL.
+Code     [Uncaptioned image] Models
 1 Introduction
-Figure 1: Comparative performance of LLDS and baseline methods on benchmark datasets. All baselines are built upon Qwen2.5-7B-Instruct. See Sec.˜6.1 for details.
-Large language models (LLMs) increasingly leverage external tools, such as search engines (jin2025search) and code-execution environments (feng2025retool; zeng2025simplerl), to augment their reasoning capabilities.
-This tool-integrated reasoning (TIR) paradigm has driven recent progress across factual question answering (jin2025search), image-based reasoning (jiang2025verltool), and mathematical problem solving (feng2025retool; jiang2025verltool).
-By enabling models to iteratively query tools, verify intermediate steps, and refine their thoughts, tool calls substantially elevate reasoning quality (qin2023toolllm).
-These advances naturally motivate the use of reinforcement learning (RL) to train LLMs to plan, interact with tools, and master multi-step decision making, as instantiated by recent GRPO-based frameworks such as Search-R1 (jin2025search).
-However, extending TIR to the RL setting introduces severe and persistent training instability.
-In particular, GRPO training in Search-R1-style jin2025search, search-integrated QA pipelines frequently exhibit sudden reward drops and catastrophic collapse (jin2025search; sun2025zerosearch).
-These failures are especially pronounced in multi-turn settings (xue2025simpletir), where tool feedback becomes part of the model’s conditioning context, creating long, highly entangled trajectories.
-Although prior work has observed these failures, the underlying mechanisms remain poorly understood.
-In this work, we firstly identify Lazy Likelihood Displacement (LLD) (deng2025effect; razin2024unintentional; ren2024learning), the stagnation or reduction of likelihood for both correct and incorrect responses during GRPO optimization (deng2025effect), as the fundamental but previously overlooked source of collapse in Search-R1–style tool-integrated reinforcement learning (TIRL).
-Using search-integrated question answering as a case study (see Fig.˜2), we show that LLD emerges early and persistently: even as rewards increase, the likelihood of correct responses enters a monotonic decline.
+Figure 1: Comparative performance of LLDS and baseline methods on benchmark datasets. All baselines are built upon Qwen2.5-7B-Instruct. See Section˜6.1 for details.
+Large language models (LLMs) increasingly leverage external tools, such as search engines (Jin et al., 2025) and code-execution environments (Feng et al., 2025; Zeng et al., 2025), to augment their reasoning capabilities.
+This tool-integrated reasoning (TIR) paradigm has driven recent progress across factual question answering (QA) (Jin et al., 2025), image-based reasoning (Jiang et al., 2025), and mathematical problem solving (Feng et al., 2025; Jiang et al., 2025).
+By enabling models to iteratively query tools, tool calls substantially elevate reasoning quality (Qin et al., 2023).
+These advances naturally motivate the use of reinforcement learning (RL) to train LLMs to plan, interact with tools, and master multi-step decision making, as instantiated by TIRL frameworks such as Search-R1 (Jin et al., 2025)
+However, extending TIR to the RL setting leads to severe and persistent training instabilities.
+In particular, GRPO training in Search-R1–style pipelines frequently suffers from abrupt reward drops and catastrophic collapse (Jin et al., 2025; Sun et al., 2025).
+These failures are especially pronounced in multi-turn tool scenarios (Xue et al., 2025), where out-of-distribution tool feedback is incorporated into the model’s context.
+Although prior work observed these failures, the underlying mechanisms lack understanding.
+In this work, we first identify Lazy Likelihood Displacement (LLD) (Deng et al., 2025; Razin et al., 2024; Ren and Sutherland, 2024), the stagnation or reduction of likelihood for both correct and incorrect responses during GRPO optimization (Deng et al., 2025), as the fundamental but previously overlooked source of collapse in TIRL.
+Using search-integrated QA as a case study (see Figure˜2), we show that LLD emerges early and persistently: even as rewards increase, the likelihood of correct responses enters a monotonic decline.
 This behavior appears across model scales, indicating that LLD is a structural failure mode of GRPO-based TIRL rather than a configuration-specific artifact.
 We further demonstrate that LLD drives a self-reinforcing LLD death spiral, where the resulting low-confidence regime amplifies negative-gradient influence from incorrect trajectories, accelerating likelihood decay, triggering entropy spikes, inflating likelihood ratios, and ultimately causing the large-gradient instability that leads to collapse.
-To counteract this failure mode, we propose a simple yet effective likelihood-preserving regularization $\rm LLDS$ that prevents harmful likelihood reductions.
-Our method integrates seamlessly with GRPO and introduces two layers of selectivity: (i) response-level gating, which activates the regularization only when a trajectory’s overall likelihood decreases, and (ii) token-level selectivity, which penalizes only the tokens responsible for the decrease.
+To counteract this failure mode, we propose a simple yet effective LLD suppression ($\rm LLDS$) regularization that prevents harmful likelihood reductions.
+Our method integrates seamlessly with GRPO and introduces two layers of selectivity: (i) action-level gating, which activates the regularization only when a response action’s overall likelihood decreases, and (ii) token-level selectivity, which penalizes only the tokens responsible for the decrease.
 This fine-grained design directly mitigates LLD while minimally interfering with GRPO’s optimization behavior.
-By preventing unintentional downward likelihood drift, $\rm LLDS$ keeps optimization away from the unstable high-gradient regime and maintains healthy training dynamics.
-Empirically, this translates into stable training, suppressed gradient explosion, and consistent gains across seven open-domain and multi-hop QA benchmarks.
+By preventing unintentional downward likelihood drift, $\rm LLDS$ maintains stable training, suppressed gradient explosion, and consistent gains across seven open-domain and multi-hop QA benchmarks.
 Our main contributions are threefold:
-  • We conduct the first systematic investigation of LLD in GRPO-driven TIRL and show that it arises with striking frequency, consistently progressing through a characteristic failure trajectory: steady likelihood decay, followed by gradient amplification and entropy explosion, and ultimately collapse.
-  • We propose a lightweight likelihood-preserving regularization $\rm LLDS$ that selectively regularize likelihood reductions and resolved the collapse issue in GRPO training.
-  • By stabilizing training, we realize significant performance gains across seven QA benchmarks (see Fig.˜1), demonstrating a more robust and reliable approach to RL-driven TIR.
-Figure 2: We illustrate the likelihood displacement in tool-integrated RL training.
-The steady-decay phase (60-120) emerges when the reward begins to increase only gradually.
-In the subsequent acceleration phase (after step 120), the likelihood of correct responses drops sharply, accompanied by a sudden surge in gradient magnitude (red star), leading to gradient explosion.
-A zoomed-in view of the acceleration region further highlights this effect, showing a clearer likelihood displacement, where the gradient accelerates rapidly while the reward starts to decline.
+$\bullet$ We identify LLD as a previously unrecognized failure mode in GRPO-based tool-integrated reinforcement learning, and show that it arises pervasively and follows a consistent, self-reinforcing trajectory of sustained likelihood decay, gradient amplification, entropy explosion, and eventual training collapse.
+$\bullet$ We propose a lightweight regularization $\rm LLDS$ that selectively regularize likelihood reductions and resolves the collapse issue in GRPO training.
+$\bullet$ By stabilizing training, we realize significant performance gains across QA benchmarks (see Figure˜1), demonstrating a more robust and reliable approach to TIRL.
 2 Related Work
 Tool-Integrated Reasoning and Agentic LLMs.
-Tool use has emerged as a powerful paradigm for equipping LLMs with adaptive reasoning capabilities.
-Early approaches relied on prompt-based orchestration (lu2023chameleon; shen2023hugginggpt) or multi-agent delegation frameworks to invoke tools without explicit training.
-Instruction-tuned models (gou2023tora; qin2023toolllm) later introduced structured tool-calling behaviors via supervised learning, but these systems remained largely static and were constrained to single-turn interactions.
-More recent work demonstrates that reinforcement learning can substantially enhance tool integration by enabling models to learn tool-usage policies through environmental feedback and task success.
-Notable systems such as RETool (feng2025retool), VERL-Tool (jiang2025verltool), and agentic LLM frameworks (maiagentic) support multi-step reasoning through dynamic tool use, self-verification, and error correction.
-This transition from static instruction-following to feedback-driven optimization has proven effective across a range of domains, including mathematical problem solving with code execution (xue2025simpletir), open-domain question answering with retrieval (jin2025search), SQL generation from natural language (jiang2025verltool), and multi-modal visual reasoning (gao2024multi).
+Tool use enables LLMs to perform adaptive, multi-step reasoning.
+Early methods relied on prompt-based (Lu et al., 2023; Shen et al., 2023) or supervised tool calling (Gou et al., 2023; Qin et al., 2023), while recent RL-based systems (e.g., RETool (Feng et al., 2025), VERL-Tool (Jiang et al., 2025)) learn tool-usage policies from environmental feedback, achieving strong performance across QA (Jin et al., 2025), code-based math reasoning (Xue et al., 2025), SQL generation (Jiang et al., 2025), and multimodal tasks (Gao et al., 2024).
 Training Collapse in Tool-Integrated GRPO.
-GRPO (guo2025deepseek) has gained popularity in reinforcement learning due to its value-free, outcome-driven formulation.
-However, applying GRPO to train LLMs for multi-turn tool-integrated reasoning remains highly unstable (jin2025search).
-When initialized from base models using only verifiable rewards (Zero RL), GRPO-trained models often experience catastrophic failure, marked by abrupt reward collapse (xue2025simpletir; maiagentic; jin2025search).
-In contrast, Proximal Policy Optimization (PPO) (schulman2017proximal) tends to exhibit more stable behavior under comparable settings (jin2025search).
-These collapse dynamics have been consistently observed across various systems, including Search-R1 (jin2025search), SimpleTIR (xue2025simpletir), and ZeroSearch (sun2025zerosearch).
-While prior work has largely focused on empirical observations of failure, only xue2025simpletir attempt to explain the phenomenon, attributing it to low-likelihood incorrect responses, which inflate importance weights and destabilize learning.
-Yet this explanation does not fully account for the structural causes of such low-likelihood trajectories.
-Notably, we observe that reward degradation often precedes the final collapse, suggesting deeper optimization issues.
-In this work, we identify Lazy Likelihood Displacement (LLD) as the core mechanism underlying GRPO’s failure mode in multi-turn TIRL.
-3 Priliminary
-TIRL enables an LLM to interact with the external tools.
-We denote the query by $\bm{x}$, the tool feedback by $\bm{o}$, and the model’s action by $\bm{y}$.
-For a given query $\bm{x}$, the model autoregressively generates an action at turn $t$ according to
+GRPO (Guo et al., 2025) is a value-free, outcome-driven RL method but exhibits severe instability in multi-turn tool-integrated settings, often collapsing when trained from base models with only verifiable rewards (Jin et al., 2025).
+This behavior is consistently observed in systems such as Search-R1 (Jin et al., 2025), SimpleTIR (Xue et al., 2025), and ZeroSearch (Sun et al., 2025), whereas PPO remains comparatively stable.
+Prior explanations attribute the collapse to training–inference mismatch (Liu et al., 2025) or low-likelihood incorrect responses (Xue et al., 2025), but do not explain PPO’s robustness or the structural origin of the instability.
+We identify Lazy Likelihood Displacement (LLD) as the underlying mechanism driving collapse in multi-turn TIRL (Section B.1).
+Recent works mitigate this via turn-level reward shaping, either relying on ground-truth documents (Zheng et al., 2025b) or an external LLM judge (Zhang et al., 2025), both introducing additional supervision or system cost and increased vulnerability to reward hacking (Guo et al., 2025).
+Recently, TreeGRPO (Ji et al., 2025) provides dense reward using tree-based rollout, but increases rollout and computational overhead.
+In contrast, our approach shows that outcome rewards alone suffice to achieve superior performance once GRPO is stabilized.
+Detailed related work see Section˜B.1.
+3 Preliminary
+We denote the query by $\bm{x}$, the tool feedback by $\bm{o}$, and the model’s action by $\bm{y}$. For a given query $\bm{x}$, the model autoregressively generates an action at turn $t$ according to
 $\pi_{\theta}\!\left(\bm{y}_{t}\mid\bm{x},\bm{y}_{0},\bm{o}_{0},\bm{y}_{1},\ldots,\bm{o}_{t-1}\right),$
@@
 If no tool is invoked, $\bm{o}_{t-1}$ is an empty string.
-The multi-turn setting therefore corresponds to trajectories that involve multiple tool calls and, consequently, multiple rounds of tool feedback.
-However, tool feedback $\mathbf{o}_{t}$ is inherently out-of-distribution (OOD) for the pretrained language model, since such feedback originates from external environments rather than the model’s own generative distribution.
-Directly forcing the model to fit these feedback tokens introduces a large training–inference mismatch and may cause the model to memorize environment-specific feedback.
-Therefore, previous works commonly mask out feedback tokens during supervised training to stabilize optimization and avoid overfitting to external content (jin2025search).
+The multi-turn setting thus corresponds to trajectories involving multiple tool calls.
+However, tool feedback $\mathbf{o}_{t}$ is inherently out-of-distribution (OOD) for the LLM, as it originates from external environments rather than the model’s generative distribution.
+Thus, prior work (Jin et al., 2025) masks feedback tokens during training to stabilize optimization.
 Tool-Integrated GRPO with Feedback Mask.
-The GRPO loss, introduced in DeepSeek-Math (shao2024deepseekmath) and DeepSeek-R1 (guo2025deepseek), enhances policy optimization by refining how reward and loss signals are computed.
-Unlike Proximal Policy Optimization (PPO) (schulman2017proximal), GRPO removes the need for value function estimation and instead uses group-relative reward normalization for optimization.
+GRPO, introduced in DeepSeek-Math (Shao et al., 2024) and DeepSeek-R1 (Guo et al., 2025) and later applied to tool use (Jin et al., 2025), is a value-free policy optimization method with group-relative reward normalization.
 Specifically, for a query–answer pair $(\bm{x},\bm{a})$, the policy $\pi_{\theta}$ samples $G$ responses:
@@
 The tool-integrated GRPO objective with feedback mask is given by:
-$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\mathbb{E}_{\begin{subarray}{c}(\bm{x},\bm{a})\sim\mathcal{D}\\ \{\bm{y}_{i,0},\ldots,\bm{y}_{i,t},\bm{o}_{i,t},\bm{y}_{i,T_{i}}\}\sim(\pi_{\theta_{\mathrm{old}},\mathcal{T}})\end{subarray}}$
-$\Bigg[\frac{1}{\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}|\hat{\bm{y}}_{i,t}|}\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}\sum_{k=1}^{|\hat{\bm{y}}_{i,t}|}\min\!\left(\gamma_{i,t,k}(\theta)\,\hat{A}_{i,t,k},\;\hat{A}_{i,t,k}\,\mathrm{clip}\big(\gamma_{i,t,k}(\theta),\,1-\varepsilon,\,1+\varepsilon\big)\right)\Bigg],$ (1)
-where $\varepsilon$ is the clipping parameter, the likelihood ratio is defined as $\gamma_{i,k}(\theta)=\frac{\pi_{\theta}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}{\pi_{\theta_{\mathrm{old}}}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}$.
-Although feedback tokens $\bm{o}$ are masked out during loss computation, they still influence the context of subsequent token predictions.
+$\mathcal{J}_{\mathrm{GRPO}}(\theta)=\mathbb{E}_{\begin{subarray}{c}(\bm{x},\bm{a})\sim\mathcal{D}\\ \{(\bm{y}_{i,t},\bm{o}_{i,t})_{t=0}^{T_{i}-1},\bm{y}_{i,T_{i}}\}\sim(\pi_{\theta_{\mathrm{old}},\mathcal{T}})\end{subarray}}$
+$\Bigg[\frac{1}{\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}|\hat{\bm{y}}_{i,t}|}\sum_{i=1}^{G}\sum_{t=1}^{T_{i}}\sum_{k=1}^{|\hat{\bm{y}}_{i,t}|}\min\!\Bigl(\gamma_{i,t,k}(\theta)\,\hat{A}_{i,t,k},\;$
+$\hat{A}_{i,t,k}\,\mathrm{clip}\big(\gamma_{i,t,k}(\theta),\,1-\varepsilon,\,1+\varepsilon\big)\Bigr)\Bigg],$ (1)
+where $\varepsilon$ is the clipping parameter, the likelihood ratio is defined as $\gamma_{i,t,k}(\theta)=\frac{\pi_{\theta}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}{\pi_{\theta_{\mathrm{old}}}(\bm{y}_{i,t,k}\mid\bm{x},\ldots,\bm{y}_{i,t-1},\bm{o}_{i,t-1},\bm{y}_{i,t,<k})}$.
+Although feedback tokens $\bm{o}$ are masked out, they still influence the context of subsequent token predictions.
+Figure 2: We illustrate the likelihood displacement in tool-integrated RL training.
+The steady-decay phase (60-120) emerges even when the reward increases.
+In the subsequent acceleration phase (after step 120), the likelihood of correct responses drops sharply, accompanied by a sudden surge in gradient magnitude (red star), leading to gradient explosion.
+A zoomed-in view of the acceleration region further highlights this effect, showing a clearer likelihood displacement, where the gradient accelerates rapidly while the reward starts to decline.
 4 Lazy Likelihood Displacement in Tool-integrated GRPO
-Recent work (deng2025effect) introduces Lazy Likelihood Displacement (LLD) in GRPO for text-based, non-tool settings, showing that the likelihood of correct responses often decreases or improves only marginally during optimization.
-In this section, we extend LLD to the tool-integrated RL regime and define:
-Definition 4.1 ( Tool-Lazy Likelihood Displacement)
-Let $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{fin}}}$ denote the initial and finetuned policies obtained before and after optimizing a preference-learning objective $\mathcal{J}$ (e.g., Sec.˜3) over a dataset $\mathcal{D}$, with $\mathcal{J}(\theta_{\mathrm{fin}})<\mathcal{J}(\theta_{\mathrm{old}})$.
+We identify a critical but previously unrecognized failure mode in tool-integrated GRPO, in which likelihood displacement progressively accelerates through compounding low-confidence responses and ultimately destabilizes training.
+While (Deng et al., 2025) identifies LLD as a training pathology in single-turn GRPO, their analysis is restricted to closed-loop text generation without external observations.
+In contrast, tool-integrated reinforcement learning fundamentally changes the learning dynamics: LLD becomes a trajectory-level instability that interacts with out-of-distribution tool feedback and prefix-shared decision branches, leading to systematic likelihood decay, gradient amplification, and eventual training collapse.
+We further show that this phenomenon arises from two structural factors unique to the tool-integrated GRPO setting (analyzed in Section 5.4) and consistently leads to catastrophic collapse behavior.
+We formalize and define LLD in the tool-integrated RL regime in this section.
+Definition 4.1 (Tool-LLD).
+Let $\pi_{\theta_{\mathrm{old}}}$ and $\pi_{\theta_{\mathrm{fin}}}$ denote the initial and finetuned policies obtained before and after optimizing the objective $\mathcal{J}$ (e.g., Section˜3) over a dataset $\mathcal{D}$, with $\mathcal{J}(\theta_{\mathrm{fin}})<\mathcal{J}(\theta_{\mathrm{old}})$.
 Consider a tool-integrated trajectory consisting of alternating actions and feedback, $(\bm{y}_{0},\bm{o}_{0},\;\bm{y}_{1},\bm{o}_{1},\;\ldots,\;\bm{y}_{T}),$ where only the actions $\{\bm{y}_{t}\}_{t=0}^{T}$ are used in likelihood computation (feedback is masked).
 For each response action $\bm{y}_{t}$, define its log-likelihood change as
-$\Delta_{t}(\mathbf{x},\bm{y}_{t}):=\ln\pi_{\theta_{\mathrm{fin}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t})-\ln\pi_{\theta_{\mathrm{old}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t}).$
-We say that Lazy Likelihood Displacement (LLD) occurs for the action $t$ if $\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon$, where $\epsilon$ is a small or non-positive constant.
-And LLD occurs for the whole response if
+$\Delta_{t}(\mathbf{x},\bm{y}_{t})$ $:=\ln\pi_{\theta_{\mathrm{fin}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t})$
+$-\ln\pi_{\theta_{\mathrm{old}}}(\bm{y}_{t}\mid\mathbf{x},\bm{y}_{<t},\bm{o}_{<t}).$
+We say that LLD occurs for the action $t$ if $\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon$, where $\epsilon$ is a small or non-positive constant. And LLD occurs for the whole response if
 $\sum_{t=0}^{T}\Delta_{t}(\mathbf{x},\bm{y}_{t})\;\leq\;\epsilon,$ (2)
-Thus, LLD captures the failure mode where one or more response actions exhibit negligible or even negative improvement in likelihood under the optimized policy.
-To better understand the origin of LLD, we highlight how negative gradients suppress the likelihood of correct actions.
-The following informal theorem (the formal version and proof see App.˜A) restates Theorem 4.4 from deng2025effect in the setting of tool-integrated RL:
-Theorem 4.2 (Informal: Trajectory-Level LLD in Tool-Integrated GRPO)
-In tool-integrated GRPO, the likelihood of an otherwise correct response can decrease when (i) incorrect responses of low likelihood and (ii) incorrect responses whose embeddings closely resemble the correct one induce large negative gradients that dominate the positive updates.
-These forces jointly lead to trajectory-level Lazy Likelihood Displacement.
-In the setting of tool-integrated GRPO, we observed with striking frequency that the likelihood of correct responses fails to improve.
-This indicates an especially acute form of LLD ($\epsilon\leq 0$), causing a progressive decay in the model’s overall output likelihood.
-This compounding decay is a distinctive failure mode of Search-R1 style tool-integrated RL.
-5 Lazy Likelihood Displacement in tool-integrated GRPO
-In this section, we illustrate the prevalence of LLD in tool-integrated GRPO training and show how the LLD ($\epsilon\leq 0$) can catastrophically drive the model into training collapse.
-5.1 Likelihood Dynamic
-To illustrate the prevalence and progression of LLD during training, we visualize the likelihood-displacement trajectory in tool-integrated RL.
-As shown in Fig.˜2, the evolution unfolds in three characteristic phases.
-Phase I (early stagnation): the likelihood of correct responses remains nearly unchanged even as the reward increases, showing the initial emergence of LLD.
-The later two phases correspond to the $\epsilon\leq 0$ regime, where the likelihood of correct responses strictly decays rather than improves.
-Phase II (steady-decay, steps 60–120): the likelihood exhibits a slow but consistent downward drift, while the reward grows only gradually and the gradient norm stays stable—indicating persistent LLD.
-Phase III (acceleration, after step 120): the likelihood begins to collapse sharply, coinciding with a rapid surge in gradient magnitude (marked by the red star), which triggers gradient explosion and ultimately leads to training collapse.
-The zoomed-in view on the right highlights this transition: likelihood keeps decreasing while gradients accelerate upward, and the reward curve begins to fall.
-Although the final collapse is caused by exploding gradients, the likelihood decay is present throughout training, introducing an accumulating instability that becomes dramatically amplified during the acceleration phase (see Appendix Sec.˜B.1 for more discussion).
-5.2 Lazy Likelihood Displacement Death Spiral
-In Fig.˜2, we observe an accelerated decline in response likelihood as training progresses.
-In this section, we first simplify notation by writing the feedback-masked trajectory as $\hat{\bm{y}}=(\bm{y}_{0},\bm{y}_{1},\ldots,\bm{y}_{T})$.
-We then formally characterize this accelerated likelihood decay as a Lazy Likelihood-Displacement (LLD) death spiral.
-Definition 5.1 (LLD Death Spiral)
-Consider a tuple $(\bm{x},\hat{\bm{y}}^{+})$ and a policy update from $\pi_{\theta_{t}}$ to $\pi_{\theta_{t+1}}$.
-We say the system enters an LLD Death Spiral if the policy evolution exhibits the following self-reinforcing progression:
-$\mathrm{LLD}_{t}\;\Longrightarrow\;C^{\mathrm{low}}_{t}\;\Longrightarrow\;\mathrm{LLD}_{t+1},\quad\epsilon_{t+1}<\epsilon_{t}\leq 0$
-where $\mathrm{LLD}_{t}$ denotes lazy likelihood displacement at time $t$ with decreasing likelihood ($\epsilon_{t}\leq 0$), and $C^{\mathrm{low}}_{t}$ denotes low-confidence trajectories whose likelihood decreases under $\pi_{\theta_{t}}$.
-The transition $\mathrm{LLD}_{t}\Longrightarrow\ C^{\mathrm{low}}_{t}$ reflects that degraded likelihood leads to increasingly diffuse predictions, while the transition $C^{\mathrm{low}}_{t}\Longrightarrow\ \mathrm{LLD}_{t+1}$ occurs because low-confidence incorrect responses often contain actions similar to those in correct responses, inducing a larger value in Eq.˜9 and thus a stronger negative-gradient influence.
-When these effects compound across iterations, the likelihood decay accelerates, creating a self-perpetuating collapse that we refer to as the LLD Death Spiral.
-Then we demonstrate the LLD Death Spiral through accelerated entropy explosion and accelerated per-sample LLD.
-Figure 3: Effect of likelihood displacement across different training iterations for the Qwen2.5-3B-Instruct model.
+Therefore, LLD accounts for the failure scenario where even correct response actions fail to boost confidence (or even reduce it) when such an enhancement is explicitly required by the loss function.
+(a) Likelihood displacement across training steps.
+(b) Mechanism
+Figure 3: (a) Effect of likelihood displacement across different training iterations for the Qwen2.5-3B-Instruct model.
 Results are computed on the first 50 samples of the training set, discarding cases where all responses are uniformly correct or uniformly incorrect.
 Bars below zero (orange) indicate samples whose correct responses’ likelihood decreases after training.
+(b) Illustration of the mechanism: as training progresses, the likelihood of correct responses decreases while hidden-state similarity increases, leading to stronger LLDS.
+We generalize the negative-gradient characterization of group-relative updates in (Deng et al., 2025) to a multi-turn setting with external observations, where the learning dynamics differ qualitatively.
+We present the following informal theorem (Formal statement and proof in Appendix˜A).
+Theorem 4.2 (Informal: Trajectory-Level LLD in Tool-Integrated GRPO).
+In tool-integrated GRPO, the likelihood of a correct response action can decrease when incorrect responses satisfy two conditions: (i) Low likelihood: The model assigns low probability to incorrect responses, causing large prediction-error weights that magnify their influence, and (ii) High embedding similarity: Incorrect responses share similar representations with correct ones, causing negative gradients to interfere with positive updates.
+As they frequently do in tool-integrated settings (see Section 5.4), negative gradients dominate, causing LLD.
+In the setting of tool-integrated GRPO, we observed with striking frequency that the likelihood of correct responses reduces (Figure˜2).
+This indicates an especially acute form of LLD ($\epsilon\leq 0$), causing a progressive decay in the model’s output likelihood.
+This compounding decay is a distinctive failure mode of Search-R1 style TIRL.
+5 LLD in Tool-Integrated GRPO Collapse
+This section examines the prevalence of LLD in tool-integrated GRPO and illustrates how sustained likelihood decay ($\epsilon\leq 0$) leads to catastrophic training collapse.
+All analyses in this section use the NQ dataset (Kwiatkowski et al., 2019) unless otherwise stated.
+5.1 Likelihood Dynamic
+To characterize the prevalence and evolution of LLD during training, we visualize the training trajectory in Figure˜2.
+The dynamics exhibit three distinct phases.
+Phase I (early stagnation).
+The likelihood of correct responses remains nearly constant despite increasing rewards, indicating the onset of LLD.
+The subsequent phases correspond to the $\epsilon\leq 0$ regime, where likelihood strictly decreases.
+Phase II (steady decay).
+Likelihood declines slowly but consistently, while rewards continue to rise and gradient norms remain stable (e.g., steps 60–120 in Figure˜2).
+Phase III (acceleration).
+After a turning point (around step 120 in Figure˜2), likelihood drops sharply and coincides with a rapid increase in gradient magnitude (marked by the red star), leading to gradient explosion and eventual training collapse.
+We further provide a zoomed-in view showing that, although collapse is ultimately triggered by gradient explosion, instability and reward degradation emerge earlier.
+5.2 LLD Death Spiral
+In Figure˜2, we observe an accelerated decline in response likelihood as training progresses. We formalize the resulting accelerated likelihood decay as an LLD death spiral.
+Definition 5.1 (Informal LLD Death Spiral).
+Consider a policy update from $\pi_{\theta_{s}}$ to $\pi_{\theta_{s+1}}$. We say the system enters an LLD Death Spiral if the policy evolution follows the self-reinforcing pattern
+$\mathrm{LLD}_{s}\;\Longrightarrow\;C_{s}^{\mathrm{low}}\;\Longrightarrow\;\mathrm{LLD}_{s+1},\qquad\varepsilon_{s+1}<\varepsilon_{s}\leq 0,$
+where $\mathrm{LLD}_{s}$ denotes lazy likelihood displacement at iteration $s$, and $C_{s}^{\mathrm{low}}$ denotes low-confidence trajectories whose likelihood decreases under $\pi_{\theta_{s}}$.
+A formal definition is provided in Section˜A.2, and the mechanism is visualized in Figure˜3.
+Beyond the acceleration phase shown in Figure˜2, we further present additional empirical evidence of the LLD death spiral in the following section.
+5.3 Empirical Evidence for LLD Death Spiral
+We demonstrate the LLD Death Spiral through accelerated entropy explosion and accelerated per-sample LLD.
 Accelerated Per-Sample LLD.
-We conduct a controlled experiment on the searching-based question answering NQ (kwiatkowski2019natural) dataset to examine how the negative gradient of GRPO affects the log-likelihood of correct responses.
-Using Qwen2.5-3B-Ins (yang2024qwen2), we generate 8 rollouts per question and retain only those examples that contain a mixture of correct and incorrect responses, discarding cases where all responses are uniformly correct or incorrect.
-To isolate the per-sample learning dynamics, we reinitialize the model parameters $\theta$ for each individual question, apply a single GRPO update to obtain $\theta^{\prime}$, and then measure the average log-likelihood change of the correct responses:
-$\Delta(\bm{x}):=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}\left[\ln\pi_{\theta^{\prime}}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})-\ln\pi_{\theta}(\hat{\bm{y}}_{i}^{+}\mid\bm{x})\right],$ (3)
-for a question $\bm{x}$ with $N^{+}$ correct responses.
-We present the results in Fig.˜3.
-In the early stage of training (iteration 50), only mild LLD is observed: most questions show negligible reductions in the likelihood of correct responses, likely because correct and incorrect trajectories remain structurally dissimilar.
-However, as training progresses, the effect becomes more pronounced.
-Increasing structural similarity between incorrect and correct responses leads to broader likelihood degradation, as shown by the rising orange curve.
-As training enters the acceleration phase (iterations 120–140), the trend intensifies into an LD death spiral, with over half of the samples at iteration 140 showing substantial drops in the likelihood of correct responses.
-This collapse is driven by the extremely low likelihood of incorrect responses, which amplifies negative-gradient contributions and destabilizes learning.
+We conduct a controlled study on the NQ dataset (Kwiatkowski et al., 2019) to examine how GRPO training affect the likelihood of correct responses.
+Using Qwen2.5-3B-Ins (Yang et al., 2024), we generate eight rollouts per question and retain only samples containing both correct and incorrect responses.
+For each question, we reinitialize model parameters, apply a single GRPO update, and measure the average log-likelihood change of correct responses:
+$\Delta(\bm{x}):=\frac{1}{N^{+}}\sum_{i=1}^{N^{+}}\sum_{t=0}^{T_{i}}\left[\ln\pi_{\theta^{\prime}}(\bm{y}_{i,t}^{+}\mid\bm{x})-\ln\pi_{\theta}(\bm{y}_{i,t}^{+}\mid\bm{x})\right],$ (3)
+where $N^{+}$ denotes the number of correct responses.
+As shown in Figure˜3(a), likelihood degradation emerges as training progresses (orange curve).
+During the acceleration phase (iterations 120–140), this effect intensifies into an LLD death spiral, with >> 50% samples exhibiting likelihood drops.
+Figure 4: We visualize the evolution of entropy, response length, and valid-search ratio during training.
+For both Qwen2.5-3B-Instruct and Qwen2.5-3B-Base, entropy rises sharply prior to collapse, indicating severe likelihood displacement.
 Accelerated Entropy Explosion.
-Fig.˜4 visualizes the evolution of entropy, response length, and valid-search ratio during training for Qwen2.5-3B-Ins (a) and Qwen2.5-3B-Base (b).
-Across both models, mean token entropy initially increases slowly, mirroring the early slow-LLD regime in Fig.˜2, but later transitions into a phase of sharply accelerated growth that aligns with the acceleration phase of the LLD trajectory.
-This steepening entropy slope is a clear demonstration of the death spiral: as low-confidence responses accumulate, the model assigns increasingly diffuse, low-probability distributions to its tokens, which further strengthens LD and drives the system toward instability.
-Crucially, this entropy acceleration occurs while response length and valid-search counts remain nearly unchanged, confirming that the rising entropy and the corresponding likelihood decay is not caused by trajectory length or tool-query frequency, but is instead a direct manifestation of the LD effect itself.
-Figure 4: We illustrate how entropy, response length, and valid-search ratio evolve during training.
-For both Qwen2.5-3B-Instruct (a) and Qwen2.5-3B-Base (b), entropy exhibits a accerlerrated upward trend prior to collapse indicating a strong LD issue.
-Meanwhile, the response length and valid-search times remain stable in the early stages but later begin to fluctuate markedly and eventually drop sharply.
-5.3 Watch for correct actions embedded within incorrect responses.
-Observing tool-integrated GRPO is substantially more susceptible to LLD than non-tool settings (deng2025effect).
-This section examines LLD at the granularity of response actions, where each action is defined by the segments between consecutive tool calls.
-Our analysis on Qwen2.5-3B-Ins (yang2024qwen2) reveals a striking pattern unique to tool-integrated GRPO: correct actions frequently appear within otherwise incorrect responses, and this intermixed structure disrupts the likelihood estimation process, amplifying LLD and destabilizing training.
-Frequent correct actions in Incorrect responses.
-In our experiments, we found that the model typically uses its first response action to generate a search query, and the correctness of this initial action increases steadily as training progresses.
-To be specific, we measure whether the retrieved documents produced by an incorrect response match those retrieved by a correct response; if so, we treat the first action as correct.
-The accuracy of this initial action across different training stages is shown in Fig.
-5.
-As indicated by the dashed green line, accuracy is low in the early stages, reflecting the model’s initial acquisition of search capability, but rises markedly over time.
-By step 140, roughly 60% of incorrect responses begin with a correct search query, highlighting the high structural similarity shared by first-action actions.
+Figure˜4 shows the evolution of entropy, response length, and valid-search ratio for Qwen2.5-3B-Ins and Qwen2.5-3B-Base.
+For both models, token entropy increases slowly in early training, consistent with the slow-LLD regime in Figure˜2, and then accelerates sharply during the LLD acceleration phase.
+This entropy surge reflects the LLD death spiral: accumulating low-confidence responses induce increasingly diffuse token distributions, further amplifying likelihood decay and instability.
+Notably, response length and valid-search frequency remain nearly constant, indicating that the entropy growth and likelihood degradation are driven by LLD rather than longer trajectories or increased tool usage.
+5.4 Why Tool-Integrated (TI) GRPO Amplifies LLD
 Figure 5: Evolution of token log-likelihood (measured before vs.
@@
 With training, both likelihoods drop while the overlap between tool observations in incorrect and correct trajectories increases, suggesting many incorrect responses begin with a correct search, which skews likelihood estimates and contributes to LLD.
-Faster likelihood decay of the first-action response.
-As illustrated in Fig.
-5, we further observe that the likelihood of the first action (light blue) declines substantially more than that of the second action (blue).
-Early in training, the first-action likelihood is higher, largely because the second action must condition on out-of-distribution feedback.
-However, as training progresses, both actions exhibit likelihood decay, with the first action eventually dropping below the second around step 110, where the first-action correctness rate is roughly 50%, indicating strong similarity between correct and incorrect trajectories.
-After step 120, this decay accelerates sharply, marking the onset of the LLD death spiral, where low-likelihood incorrect partial signals further amplify degradation.
-Nonsensical first-action.
-As LLD intensifies and the likelihood of the first action collapses, the model begins producing nonsensical outputs.
-As illustrated in Fig.˜15 in the appendix, it eventually generates random, meaningless tokens, an effect driven by the severely reduced response likelihood, which causes sampling to select arbitrary words.
-Although full collapse may not yet have occurred at this stage, the model is already effectively unusable, as it can no longer produce meaningful responses.
-As a result, it becomes essential to reduce the unintended penalization of correct actions within incorrect responses. A more detailed discussion is provided in Sec.˜7.
-5.4 LLD regularization
-To address Lazy Likelihood Displacement (LLD), we introduce a family of likelihood-preserving regularizers that prevent the model from unintentionally reducing the likelihood of responses during GRPO training.
-Specifically, for the given preserving responses $\bm{y}_{i}\in\mathcal{Y}_{\rm pre}$, we compare its token-level likelihood under the old policy (previous step) and the fine-tuned policy.
-Our base regularizer penalizes only the tokens whose likelihood decreases after an update.
-LLD: Token-Level Likelihood Preservation. LLD applies a token-level penalty: within each response $\mathbf{y}_{i}$, only likelihood-reducing tokens contribute to the loss:
-$L_{\rm LLD}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (4)
-This token-level selectivity ensures that the model is penalized only for genuinely harmful likelihood reductions, without interfering with updates that improve the response overall.
-However, some responses may still be globally improving even if a few individual tokens decrease; penalizing such cases may introduce overly strong constraints.
-LLDS: Response-Level Gating.
-To avoid unnecessary penalties on globally improving responses, LLDS introduces a response-level gating mechanism: the penalty activates only when the total likelihood of the response decreases.
-The LLDS loss is:
-$L_{\rm LLDS}=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}|\mathbf{y}_{i}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}$ $\underbrace{\mathbbm{1}\!\left[\sum_{y_{i}\in\mathbf{y}_{i}}\left(\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)>0\right]}_{\text{Activated only when sum > 0}}\cdot$
-$\sum_{y_{i}\in\mathbf{y}_{i}}\underbrace{\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Likelihood-reducing tokens}}$ (5)
-This structure preserves normal GRPO learning while directly suppressing responses that suffer from LLD.
-LLDS-MA: Masking Answer Tokens.
-To further encourage multi-step reasoning and tool usage, we mask out final answer tokens from the regularization term.
-LLDS-MA penalizes likelihood reductions only on reasoning and tool-interaction tokens, excluding the answer span $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}$}:
-$\underbrace{\sum_{y_{i}\in\mathbf{y}_{i}/\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}}}\max\left(0,\;\ln\pi_{\theta_{\rm old}}(y_{i}|\mathbf{x},\mathbf{y}_{<i})-\ln\pi_{\theta}(y_{i}|\mathbf{y}_{<i})\right)}_{\text{Mask Answer Likelihood-reducing tokens}}$ (6)
-Finally, We integrate the regularization into the GRPO objective as
-$L_{\rm total}=L_{\rm GRPO}+\lambda L_{\rm{LLDS(-MA)}}$ (7)
-where $\lambda$ is the regularization weight.
-We use LLDS as the default variant and switch to LLDS-MA when stronger encouragement of tool usage is desired.
-For the preserving set $\mathbf{y}]_{\rm pre}$, we include all responses with non-negative advantages $\hat{A}\geq 0$, ensuring that correct responses ($\hat{A}>0$) and untrained responses ($\hat{A}=0$) do not suffer likelihood reduction.
-The effect of $\lambda$ is examined empirically in Fig.˜8.
-6 Experiments and Analysis
-We evaluate the empirical effectiveness of $L_{\rm LLDS(-MA)}$ through comprehensive experiments.
-Experimental settings.
-For training, we follow the setup in jin2025search and conduct experiments using two model families: Qwen-2.5-3B and Qwen-2.5-7B, each in both Base and Instruct variants (yang2024qwen2).
-We consider two training configurations.
-(1) NQ-Only (single-hop): the model is trained solely on the single-hop Natural Questions (NQ) dataset (kwiatkowski2019natural).
-(2) NQ+Hotpot (single-hop+multi-hop): the model is trained on a merged corpus combining NQ with HotpotQA (yang2018hotpotqa), providing broader coverage of both open-domain and multi-hop reasoning.
-For the retrieval dataset, we use the 2018 Wikipedia dump (karpukhin2020dense) as the knowledge corpus and employ E5 (wang2022text) as the dense retriever.
-To ensure fair comparison across retrieval-augmented baselines, we fix the number of retrieved passages to three, following the configuration in jin2025search.
-Unless otherwise specified, we use the same optimization hyperparameters as Search-R1 (jin2025search), with the only modification being a reduced maximum turn limit: from 4 in Search-R1 to 2 for NQ-only training and 3 for the NQ+Hotpot setting, improving training efficiency.
-Unless stated otherwise, the regularization weight is fixed at $\lambda=0.1$.
-Evaluation settings.
-The Evaluation is performed on the validation or test sets of seven datasets, categorized as follows: (1) General Question Answering: NQ (kwiatkowski2019natural), TriviaQA (joshi2017triviaqa), and PopQA (mallen2022not).
-(2) Multi-Hop Question Answering: HotpotQA (yang2018hotpotqa), 2WikiMultiHopQA (ho2020constructing), Musique (trivedi2022musique), and Bamboogle (press2022measuring).
-We adopt exact match (EM) as the primary evaluation metric, consistent with the protocol in jin2025search.
-Such diversity ensures a thorough examination of search engine-integrated LLM performance across a range of realistic reasoning scenarios.
-6.1 Experimental Results
+We then show that the severity of LLD in TI-GRPO stems from two interacting factors predicted by Theorem˜4.2 and demonstrated in Figure˜3(b): high response similarity between correct and incorrect trajectories, and low likelihood responses that induce large prediction-error weights.
+Frequent correct actions in incorrect responses.
+Our analysis on Qwen2.5-3B-Ins that trained on the NQ dataset reveals a striking pattern: correct actions frequently appear within otherwise incorrect responses.
+Specifically, in TI GRPO, the first response action typically formulates a search query, whose correctness increases steadily during training regardless of final answer.
+We label the first action of incorrect responses as correct if its retrieved documents match those of a correct response.
+As shown in Fig.
+5, this accuracy (green line) is initially low but rises sharply, reaching $\sim$60% by step 140, indicating high similarity on the first action between correct and incorrect trajectories.
+At the same time, the likelihood of the first action (light blue) decays faster than that of the second action (blue): although initially higher due to the second action conditioning on OOD feedback, it drops below the second action around step 110, where first-action correctness reaches $\sim$50%.
+After step 120, this decay accelerates sharply, where low-likelihood incorrect prefixes further amplify degradation.
+As LLD intensifies, the likelihood of the first action collapses, and the model begins producing random, nonsensical tokens (Appendix, Fig.
+15), rendering responses effectively unusable even before full training collapse.
+These results motivate mitigating unintended penalization of correct actions within incorrect trajectories (see Appendix˜D).
+Figure 6: Effect of off-policy tool context on negative log-likelihood. Increasing the number of tool-interaction rounds consistently reduces response likelihood.
+Low Likelihood reasoning after tool feedback.
+A second factor that amplifies LLD is the use of OOD tool feedback as a conditioning context.
+They introduce a persistent distribution mismatch between the policy and its input context, systematically reducing the likelihood assigned to subsequent actions and increasing negative log-likelihood.
+As shown in Figure˜6, computed on the first 50 NQ training samples, this effect strengthens with the number of tool-interaction rounds, producing a monotonic increase in NLL for both Qwen2.5-3B-Instruct and Qwen2.5-7B-Instruct models.
+5.5 LLD Suppression regularization
+To mitigate LLD, we propose a likelihood-preserving regularizer, $\rm LLDS$, which prevents unintended reductions in response likelihood during GRPO training.
+Specifically, for each given preserving response $\bm{y}_{i}\in\mathcal{Y}_{\rm pre}$, we compare token-level likelihoods under the previous and current policies, and quantify their change by
+$\Delta_{i,t,k}$ $=\ln\pi_{\theta_{\mathrm{old}}}\!\left(y_{i,t,k}\mid\mathbf{x},\bm{y}_{i,<t},\bm{o}_{i,<t},\bm{y}_{i,t,<k}\right)$
+$-\ln\pi_{\theta}\!\left(y_{i,t,k}\mid\mathbf{x},\bm{y}_{i,<t},\bm{o}_{i,<t},\bm{y}_{i,t,<k}\right).$
+where $i$ indexes samples, $t$ actions (turns), and $k$ tokens. Positive $\Delta_{i,t,k}$ indicates a likelihood decrease after the update. These tokens are candidates for regularization.
 Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
 NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
+Qwen2.5-3b-Base/Instruct
 Direct Inference 0.106 0.288 0.108 0.167 0.149 0.244 0.020 0.024 0.109 0.134
-CoT 0.023 0.032 0.005 0.020 0.021 0.021 0.002 0.000 0.011 0.015
-IRCoT 0.111 0.312 0.200 0.208 0.164 0.171 0.067 0.240 0.161 0.181
 Search-o1 0.238 0.472 0.262 0.324 0.221 0.218 0.054 0.320 0.203 0.255
 RAG 0.348 0.544 0.387 0.426 0.255 0.226 0.047 0.080 0.152 0.270
-SFT 0.249 0.292 0.104 0.215 0.186 0.248 0.044 0.112 0.147 0.176
 R1-base 0.226 0.455 0.173 0.285 0.201 0.268 0.055 0.224 0.187 0.229
@@
 Search-R1-PPO-Ins 0.341 0.545 0.378 0.421 0.324 0.319 0.103 0.264 0.253 0.325
+ZeroSearch-Base 0.430 0.616 0.414 0.487 0.338 0.346 0.130 0.139 0.238 0.345
+ZeroSearch-Ins 0.414 0.574 0.448 0.479 0.274 0.300 0.098 0.111 0.196 0.317
+StepSearch 0.296 0.490 0.341 0.375 0.294 0.360 0.140 0.258 0.263 0.311
+CriticSearch - - - - 0.414 0.409 0.180 0.368 0.343 -
+Tree-GRPO 0.468 0.597 0.436 0.500 0.424 0.437 0.178 0.432 0.368 0.424
 Qwen2.5-3b-Base
-NQ-Only
-Search-R1-GRPO 0.440 0.582 0.413 0.478 0.265 0.244 0.061 0.113 0.171 0.303
-+LLD 0.462 0.609 0.464 0.512 0.286 0.248 0.063 0.113 0.178 0.321 (+5.9%)
-+LLDS 0.478 0.605 0.449 0.511 0.288 0.250 0.062 0.129 0.182 0.323 (+6.6%)
-NQ+hotpot
 Search-R1-GRPO 0.421 0.583 0.413 0.472 0.297 0.274 0.066 0.128 0.191 0.312
-+LLDS 0.479 0.634 0.462 0.525 0.350 0.308 0.086 0.201 0.236 0.360 (+15.4%)
-+LLDS-MA 0.486 0.632 0.462 0.527 0.444 0.430 0.183 0.371 0.357 0.430 (+37.8%)
++LLDS 0.479 0.627 0.453 0.520 0.345 0.323 0.082 0.210 0.240 0.360
++LLDS-MA 0.483 0.633 0.474 0.530 0.452 0.443 0.197 0.395 0.372 0.440 (+45.2%)
 Qwen2.5-3b-Ins
 Search-R1-GRPO 0.397 0.565 0.391 0.451 0.331 0.310 0.124 0.232 0.249 0.336
-NQ-Only
-+LLDS 0.490 0.610 0.450 0.517 0.294 0.193 0.074 0.121 0.171 0.319
-+LLDS-MA 0.464 0.616 0.458 0.513 0.363 0.320 0.131 0.315 0.282 0.381 (+13.4%)
-NQ+hotpot
-+LLDS 0.462 0.622 0.460 0.515 0.432 0.383 0.180 0.395 0.347 0.419 (+24.7%)
++LLDS 0.462 0.627 0.468 0.519 0.443 0.447 0.197 0.444 0.383 0.441 (+31.3%)
+Search-R1-GSPO 0.376 0.561 0.374 0.437 0.332 0.329 0.098 0.250 0.252 0.331
++LLDS 0.489 0.640 0.483 0.537 0.466 0.456 0.215 0.403 0.385 0.451(+36.3%)
 Table 1: Results on General QA and Multi-Hop QA datasets for Qwen2.5-3b-Base/Instruct.
-denotes results from (jin2025search) using NQ+hotpot training.
-denotes open-domain datasets and denotes retrieval-augmented datasets.
-We evaluate our proposed LLD-mitigation strategy on seven open-domain and multi-hop QA benchmarks using two model families: Qwen2.5-3B (Base/Instruct) and Qwen2.5-7B (Base/Instruct).
-Tables 1 and 2 summarize the EM performance across all settings.
-Since vanilla GRPO training often collapses and drives the reward to zero, we use the results from Search-R1 jin2025search, corresponding to the best checkpoint before collapse, as the GRPO baseline.
-Results on Qwen2.5-3B.
-As shown in Tab.˜1, GRPO attains an average score of 0.303 in the NQ-only setting.
-Incorporating LLD raises performance to 0.321 (+5.9%), and LLDS provides a further improvement to 0.323 (+6.6%).
-In the more challenging NQ+Hotpot setting, LLDS increases the vanilla GRPO score from 0.312 to 0.360, a 15.4% relative gain.
-Notably, since LLDS tends to limit the base model to invoking search only once, we apply the multi-aspect variant LLDS-MA, which delivers the strongest performance with an average score of 0.430, representing a substantial +37.8% improvement ratio over vanilla GRPO.
-A similar trend is observed for Qwen2.5-3B-Instruct: LLDS-MA improves performance to 0.381 under NQ-only, and LLDS achieves 0.419 in the NQ+Hotpot setting (+24.7%).
+All results are reported using Exact Match (EM).
+denotes results from (Jin et al., 2025).
+denotes results from (Zhang et al., 2025) that only have multi-hop QA results.
+denotes in-distribution datasets, while denotes out-of-distribution datasets.
+$\rm LLDS$: Action-Level Gating.
+To avoid unnecessary penalties on globally improving responses, we introduce an action-level gating mechanism: the penalty activates only when the total likelihood of an action decreases.
+The resulting $\rm LLDS$ loss is defined as:
+$L_{\mathrm{LLDS}}$ $=\frac{1}{\sum_{i=1}^{|\mathcal{Y}_{\mathrm{pre}}|}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{i=1}^{|\mathcal{Y}_{\mathrm{pre}}|}\sum_{t=0}^{T_{i}}$
+$\underbrace{\mathbbm{1}\!\left[\sum_{k=1}^{|\bm{y}_{i,t}|}\Delta_{i,t,k}>0\right]}_{\text{Activated only when sum > 0}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{Likelihood-reducing tokens}},$ (4)
+This structure directly suppresses actions that suffer from LLD.
+We also introduce token-level and response-level variants in Section˜B.3, unless stated otherwise, $\rm LLDS$ refers to the action-level gating.
+LLDS-MA: Masking Answer Tokens.
+To further promote multi-turn reasoning and tool usage, we assign smaller regularization weights to final answer tokens, discouraging premature answer generation.
+Consequently, $\rm LLDS$-MA penalizes likelihood reductions on reasoning and tool-interaction tokens more strongly, while only softly regularizing the answer span $\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}$}:
+$\sum_{k=1}^{|\mathbf{y}_{i,t}|}w_{i,t,k}\;\max\!\left(0,\;\Delta_{i,t,k}\right),w_{i,t,k}=\begin{cases}\beta,&k\in{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\bm{y}_{i,\rm Ans}}\\ 1,&\text{otherwise}\end{cases}$ (5)
+Finally, we integrate the regularization into the GRPO objective as
+$L_{\rm total}=L_{\rm GRPO}+\lambda L_{\rm{LLDS(-MA)}}$ (6)
+where $\lambda$ is the regularization weight.
+We use LLDS as the default variant, and adopt LLDS-MA when the model collapses to single-turn behavior.
+The preserving set $\mathcal{Y}_{\rm pre}$ includes all responses with non-negative advantages $\hat{A}\geq 0$, ensuring that correct responses ($\hat{A}>0$) and untrained responses ($\hat{A}=0$) do not suffer LLD.
+The effect of $\lambda$ is examined empirically in Figure˜8.
 Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
@@
 Direct Inference 0.134 0.408 0.140 0.227 0.183 0.250 0.031 0.120 0.146 0.181
-CoT 0.048 0.185 0.054 0.096 0.092 0.111 0.022 0.232 0.114 0.106
-IRCoT 0.224 0.478 0.301 0.334 0.133 0.149 0.072 0.224 0.145 0.239
 Search-o1 0.151 0.443 0.131 0.242 0.187 0.176 0.062 0.296 0.180 0.206
 RAG 0.349 0.585 0.392 0.442 0.299 0.235 0.058 0.208 0.200 0.304
-SFT 0.318 0.354 0.121 0.264 0.217 0.259 0.066 0.112 0.164 0.207
 R1-base 0.297 0.539 0.199 0.345 0.242 0.273 0.083 0.203 0.200 0.262
@@
 Search-R1-PPO-Ins 0.393 0.610 0.397 0.467 0.370 0.414 0.146 0.368 0.325 0.385
+ZeroSearch-Base 0.424 0.664 0.604 0.564 0.320 0.340 0.180 0.333 0.293 0.409
+ZeroSearch-Ins 0.436 0.652 0.488 0.525 0.346 0.352 0.184 0.278 0.290 0.391
+StepSearch 0.364 0.565 0.390 0.440 0.346 0.394 0.183 0.444 0.342 0.382
+CriticSearch - - - - 0.442 0.428 0.194 0.472 0.384 -
+Tree-GRPO 0.481 0.633 0.452 0.522 0.446 0.423 0.202 0.440 0.378 0.439
 Qwen2.5-7b-Base
-Search-R1 (GRPO) 0.395 0.560 0.388 0.448 0.326 0.297 0.125 0.360 0.277 0.350
-NQ-Only
-+LLDS 0.514 0.666 0.470 0.550 0.403 0.368 0.151 0.347 0.317 0.417 (+19.1%)
-NQ+hotpot
-+LLDS 0.504 0.666 0.471 0.547 0.465 0.475 0.209 0.444 0.398 0.462 (+32.0%)
+Search-R1-GRPO 0.395 0.560 0.388 0.448 0.326 0.297 0.125 0.360 0.277 0.350
++LLDS 0.512 0.672 0.480 0.555 0.486 0.469 0.232 0.508 0.424 0.480 (+37.1%)
 Qwen2.5-7b-Ins
-Search-R1 (GRPO) 0.429 0.623 0.427 0.493 0.386 0.346 0.162 0.400 0.324 0.396
-NQ-Only
-+LLDS 0.517 0.645 0.460 0.541 0.401 0.298 0.158 0.411 0.317 0.413 (+4.3%)
-NQ+hotpot
-+LLDS 0.498 0.662 0.464 0.541 0.473 0.443 0.228 0.516 0.428 0.469 (+18.4%)
+Search-R1-GRPO 0.429 0.623 0.427 0.493 0.386 0.346 0.162 0.400 0.324 0.396
++LLDS 0.518 0.676 0.486 0.560 0.492 0.490 0.228 0.492 0.426 0.483 (+22.0%)
 Table 2: Performance comparison on General QA and Multi-Hop QA datasets for Qwen2.5-7b-Base/Instruct.
-Results on Qwen2.5-7B As shown in Tab.˜2, for Qwen2.5-7B-Base, applying LLDS yields a substantial improvement, reaching an average score of 0.417 when trained on NQ only, representing a 19.1% relative gain over the GRPO baseline trained on NQ+Hotpot.
-When LLDS is applied under the NQ+Hotpot training setting, performance increases from 0.350 to 0.462, corresponding to an even larger 32.0% improvement.
-LLDS also achieves the strongest results across nearly all individual datasets, indicating enhanced evidence retrieval and multi-step reasoning capability.
-For Qwen2.5-7B-Instruct, we observe a similar pattern.
-Under the NQ+Hotpot training setting, LLDS raises the average score from 0.396 to 0.469 (improved 18.4%).
-Notably, LLDS achieves the best performance on every Multi-Hop QA benchmark, including 0.473 on 2Wiki, 0.443 on Musique, and 0.516 on Bamboogle.
-These results highlight LLDS(-MA) as an effective method for stabilizing GRPO training and substantially enhancing multi-turn reasoning.
+6 Experiments and Analysis
+Experimental settings.
+For training, we follow the setup in (Jin et al., 2025) and conduct experiments using two model families: Qwen-2.5-3B and Qwen-2.5-7B, each in both Base and Instruct variants (Yang et al., 2024).
+We follow the training configuration in Search-R1 (Jin et al., 2025) and use NQ+Hotpot (single-hop+multi-hop): the model is trained on a merged corpus combining NQ with HotpotQA (Yang et al., 2018), providing broader coverage of both open-domain and multi-hop reasoning.
+For the retrieval dataset, we use the 2018 Wikipedia dump (Karpukhin et al., 2020) as the knowledge corpus and employ E5 (Wang et al., 2022) as the dense retriever and fix the number of retrieved passages to three.
+Unless otherwise specified, we use the same optimization hyperparameters as Search-R1 (Jin et al., 2025), with the only modification being a reduced maximum turn from 4 to 3 for the NQ+Hotpot, improving training efficiency.
+Unless stated otherwise, the regularization weight is fixed at $\lambda=0.2$.
+More details see Sec.
+B.2.1.
+Evaluation settings.
+The Evaluation is performed on seven datasets, categorized as follows: (1) General QA: NQ (Kwiatkowski et al., 2019), TriviaQA (Joshi et al., 2017), and PopQA (Mallen et al., 2022).
+(2) Multi-Hop QA: HotpotQA (Yang et al., 2018), 2WikiMultiHopQA (Ho et al., 2020), Musique (Trivedi et al., 2022), and Bamboogle (Press et al., 2022).
+We follow (Jin et al., 2025) and adopt exact match (EM) as the evaluation metric.
+Baseline Methods.
+To evaluate the effectiveness of $\rm LLDS$, we use baselines in Search-R1 (Jin et al., 2025) and additionally include methods that improve training stability via dense rewards, namely StepSearch (Zheng et al., 2025b), CriticSearch (Zhang et al., 2025), and TreeGRPO (Ji et al., 2025).
+See Section˜B.2.2 for detailed descriptions.
+Figure 7: Comparisons of training reward curves between baseline GRPO (blue) and GRPO + LLDS (red).
+From left to right: Qwen-2.5-3B-Base, Qwen-2.5-3B-Instruct, Qwen-2.5-7B-Base, and Qwen-2.5-7B-Instruct.
+The baseline consistently collapses (reward drops to zero), while LLDS stabilizes training and sustains high rewards.
+6.1 Experimental Results
+We evaluate our proposed $\rm LLDS$ on seven open-domain and multi-hop QA benchmarks using two model families: Qwen2.5-3B (Base/Instruct) and Qwen2.5-7B (Base/Instruct).
+Tables 1 and 2 summarize the EM performance across all settings.
+Since vanilla GRPO training often collapses, we use the results from Search-R1 (Jin et al., 2025), corresponding to the best checkpoint before collapse, as the GRPO baseline.
+This choice favors GRPO and avoids underestimating its performance.
+More results see Section˜B.4.
+Results on Qwen2.5-3B.
+As shown in Table˜1, $\rm LLDS$ improves the vanilla GRPO score from 0.312 to 0.360, corresponding to a 15.4% relative gain.
+Notably, because the 3B base model tends to invoke search only once, we apply $\rm LLDS$-MA, which achieves the best performance with an average score of 0.440, representing a substantial +45.2% improvement over GRPO.
+A similar trend is observed for Qwen2.5-3B-Instruct, where LLDS attains 0.441 under the NQ+Hotpot setting (+31.3%).
+Results on Qwen2.5-7B.
+As shown in Table˜2, $\rm LLDS$ delivers substantial gains over vanilla GRPO for both base and instruct variants.
+For Qwen2.5-7B-Base, it improves the average score from 0.350 to 0.480, a 37.1% relative increase, while for Qwen2.5-7B-Instruct, performance rises from 0.396 to 0.483, corresponding to a 22.0% gain.
+Notably, $\rm LLDS$ attains the best results on all Multi-Hop QA benchmarks with MH-Avg accuracy being 0.426.
+Comparison with Dense Rewards.
+As shown in Tables˜1 and 2, stabilizing outcome-reward GRPO training with $\rm LLDS$ consistently outperforms methods that rely on dense reward supervision.
+These dense-reward approaches typically require turn-level ground-truth signals, external LLM judges, or more expensive tree-based rollouts.
+In contrast, $\rm LLDS$ achieves superior performance using only outcome rewards.
+Notably, $\rm LLDS$ outperforms the strongest dense-reward baseline, Tree-GRPO, by 4% and 10% on the 3B and 7B models, respectively, and exceeds other dense-reward methods by over 30%, demonstrating both improved effectiveness and substantially lower supervision.
+[…]
 6.2 Ablation Study and Analysis
-Impact of response-level gating.
-Applying response-level gating reduces the strength of the regularization applied during training, leading to measurable improvements on multi-hop QA tasks.
-As shown in Tab.˜1, although the average performance increases modestly by 0.2%, the method yields a notable 1.6% gain on the Bamboogle dataset, highlighting its effectiveness in enhancing multi-hop reasoning.
-NQ vs.
-NQ+Hotpot.
-Models trained only on NQ achieve strong performance on open-domain, single-hop QA but show limited ability to generalize to multi-hop settings.
-Moving to NQ+Hotpot consistently improves multi-step reasoning: for example, on Qwen2.5-3B-Base, the GRPO baseline increases from 0.303 (NQ-only) to 0.312 (NQ+Hotpot), and LLDS increases from 0.323 to 0.360 under the same shift.
-A similar trend appears for Qwen2.5-7B-Base, where LLDS rises from 0.417 (NQ-only) to 0.462 (NQ+Hotpot).
-These gains show that adding multi-hop HotpotQA, whose questions require multi-turn reasoning to answer correctly and receive reward, encourages the model to retrieve, integrate, and reason over multiple pieces of evidence more effectively than training on NQ alone.
+Impact of Regularization Strength.
+We ablate $\lambda\in\{0,0.01,0.1,0.2\}$ and plot the corresponding training reward dynamics in Figure˜8.
+As shown, Vanilla GRPO ($\lambda=0$) collapses around step 200.
+A small regularization value ($\lambda=0.01$, orange curve) delays but does not prevent collapse, which occurs around step 220.
+In contrast, a stronger regularization ($\lambda=0.2$) stabilizes training entirely, allowing the model to continue smoothly without collapse.
+We truncate the plot at step 500 for clarity, although training can proceed well beyond this point.
+Figure 8: Impact of regularization strength in training Qwen2.5-3B-Instruct on NQ and HotpotQA.
 Impact of masking answer (MA).
-We apply masking answer (MA) specifically to model variants where vanilla GRPO and LLDS degenerate into issuing only a single search call, limiting their ability to collect sufficient evidence for multi-hop reasoning.
-MA disables LLDS regularization on the final answer tokens, encouraging the model to perform additional search steps or intermediate reasoning.
-As shown in Tab.˜1, adding MA on top of LLDS yields substantial improvements for these one-search–limited models.
-LLDS-MA increases the Qwen2.5-3B-Base average score in the NQ+Hotpot setting from 0.360 (LLDS) to 0.430, and similarly improves the NQ-only Instruct variant from 0.319 to 0.381.
+We apply MA only to model variants where vanilla GRPO degenerate into issuing only a single search call.
+MA reduces LLDS regularization on the final answer tokens, encouraging the model to perform additional search steps or intermediate reasoning.
+As shown in Table˜1, LLDS-MA increases the Qwen2.5-3B-Base average score in the NQ+Hotpot setting from 0.360 (LLDS) to 0.430.
 These gains demonstrate that MA effectively encourages deeper, multi-step reasoning in scenarios where the underlying GRPO policy underutilizes search actions.
-More detailed analysis see Sec.˜B.3.
-6.3 Effect of LLDS on Training Stability across Models
-To verify the universality of the training collapse issue in GRPO and the robustness of our proposed solution, we extended our experiments across different model scales and alignment stages.
-Fig.˜6 illustrates the training reward curves for Qwen-2.5 3B and 7B models, covering both Base and Instruct versions.
-As observed in the blue curves, the baseline GRPO consistently suffers from catastrophic collapse within the first 300 steps, regardless of the model size (3B vs.
-7B) or type (Base vs.
-Instruct).
-This confirms that the instability is an inherent characteristic of the algorithm rather than an artifact of a specific model configuration.
-In contrast, the integration of LLDS (red curves) effectively mitigates this issue.
-In all four scenarios, the models trained with LLDS maintain a steady upward trajectory in rewards, successfully bypassing the collapse points that affect the baseline.
-This demonstrates that LLDS serves as a generalized and effective regularizer for stabilizing GRPO training.
-(a) Qwen-2.5-3B-Base
-(b) Qwen-2.5-3B-Instruct
-(c) Qwen-2.5-7B-Base
-(d) Qwen-2.5-7B-Instruct
-Figure 6: Comparisons of training reward curves between baseline GRPO (blue) and GRPO + LLDS (red).
-The baseline consistently collapses (reward drops to zero) across all model sizes, while LLDS stabilizes training and sustains high rewards.
-7 More Discussion and Guideline
-Based on our analysis of Lazy Likelihood Displacement (LLD) and the emergence of the LLD Death Spiral, we consolidate several practical guidelines for stabilizing tool-integrated GRPO.
-Each recommendation follows directly from the core failure mechanisms identified in our study.
-Understand why tool-integrated GRPO is uniquely vulnerable to LLD.
-Unlike free-form RL on text, tool-augmented agents introduce structural conditions that magnify likelihood drift.
-First, tool calls inject inherently out-of-distribution tokens—such as search results, API outputs, or error messages—that differ sharply from pretrained language distributions.
-These OOD segments raise token-level uncertainty and make GRPO’s relative updates more volatile, accelerating the onset of likelihood displacement.
-Second, tool-based reasoning unfolds across multiple stages, with early stages (e.g., query formulation) stabilizing more quickly than later ones (e.g., tool-result interpretation).
-Because GRPO applies a single scalar reward to all tokens, early-stage tokens that share nearly identical prefixes across correct and incorrect trajectories receive conflicting gradient signals.
-This reward–token misalignment disproportionately harms stable prefixes and amplifies LLD.
-Recognizing these structural challenges is essential for diagnosing and mitigating instability in tool-integrated RL systems.
-Closely monitor likelihood dynamics—reward alone is insufficient.
-A central insight of our analysis is that likelihood degradation begins long before any visible drop in reward.
-Across models, reward continues to rise throughout both the early and steady-decay phases, even as the likelihood of correct responses is already drifting downward.
-Because GRPO updates are governed by likelihood ratios, such early degradation silently amplifies gradients and initiates instability.
-Monitoring only the reward therefore obscures the onset of LLD.
-In practice, reliable early-warning signals come from tracking action-level and trajectory-level log-likelihood, inspecting entropy trajectories, which surge sharply once likelihood begins to collapse, and watching for sudden spikes in likelihood ratios that indicate gradient amplification.
-Visualization of likelihood or entropy curves provides the earliest and most trustworthy indication that training is approaching the unstable regime.
+More detailed analysis see Section˜B.4.4.
+Effect of LLDS on Training Stability across Models To assess the universality of GRPO collapse and the robustness of our method, we evaluate across model scales (3B and 7B) and alignment stages (Base and Instruct).
+As shown in Figure˜7, vanilla GRPO consistently collapses within the first 300 steps regardless of model size or variant.
+In contrast, integrating $\rm LLDS$ (red line) reliably stabilizes training and maintains steady reward growth across all four settings, successfully avoiding the collapse issue.
+These results demonstrate that $\rm LLDS$ serves as a general and effective regularizer for stabilizing GRPO training.
 7 Conclusion
-In this work, we firstly identified Lazy Likelihood Displacement as the central cause of instability in GRPO-based, Search-R1–style tool-integrated RL.
-We showed that LLD emerges early, occurs frequently, and reliably drives training into a self-reinforcing LLD Death Spiral characterized by monotonic likelihood decay, entropy spikes, gradient inflation, and eventual collapse.
-Our analysis provides the first systematic evidence that GRPO is uniquely vulnerable to likelihood drift in tool-integrated settings.
-To counteract this failure mode, we proposed $\rm LLDS$, a simple and effective likelihood-preserving regularizer that activates only when likelihood decreases and targets only the offending tokens.
-Empirically, $\rm LLDS$ stabilizes optimization and yields consistent gains across seven QA benchmarks, with improvements of up to +37.8% on Qwen2.5-3B and +32.0% on Qwen2.5-7B.
-These findings highlight likelihood control as a key mechanism for reliable RL-based tool use and offer a promising foundation for more robust and scalable training methods in future agentic LLMs.
+In this work, we identify LLD as the primary cause of failure in GRPO-based, search-based tool-integrated reinforcement learning, and show that it emerges early and evolves into a self-reinforcing failure mode characterized by likelihood decay, entropy inflation, and gradient explosion.
+Our analysis reveals that tool integration fundamentally reshapes group-relative optimization dynamics by coupling likelihood updates across trajectories, leading to a class of instability unique to agentic, tool-using LLMs.
+To counteract this failure mode, we propose $\rm LLDS$, a simple and effective likelihood-preserving regularizer that activates only when likelihood decreases and targets only the offending tokens, thereby stabilizing training and delivering consistent gains across seven QA benchmarks.
+More broadly, we highlight likelihood dynamics as a principled signal for early failure detection, and suggest that extending $\rm LLDS$ to other tools and agentic settings is a promising direction for future work.
 Acknowledgments: This work was partially funded by the NSERC Discovery Grant RGPIN-2021-03677, Alliance Grant ALLRP 581098-22, the Natural Science and Engineering Research Council of Canada (NSERC), the Canada CIFAR AI Chairs program, the Canada Research Chair program, an IITP grant funded by MSIT, and the Digital Research Alliance of Canada.
@@
 In this section, we first give the formal version of  Theorem˜4.2:
-Theorem A.1 (Trajectory-Level LLD in Tool-Integrated GRPO)
+Theorem A.1 (Trajectory-Level LLD in Tool-Integrated GRPO).
 Consider a tool-integrated trajectory with only the response actions $\{\bm{y}_{t}\}_{t=0}^{T}$ contribute to the likelihood, and tool-feedback tokens $\bm{o}_{t}$ are masked during likelihood computation but remain in context.
@@
   2. 2.
-    Embedding similarity: When incorrect responses are similar to correct ones, their prefix representations exhibit large inner products, amplifying the negative contribution.
+    Embedding similarity: When incorrect responses are similar to correct ones, their representations exhibit large inner products, amplifying the negative contribution.
 This high representational overlap means the model struggles to disentangle correct from incorrect continuations, causing negative examples to push gradients in harmful directions and make the model unconfident.
@@
 […]
-Theorem A.2 (Action-level)
+Theorem A.2 (Action-level).
 For any $\bm{x}$, any time $s\geq 0$, and any correct response $\bm{y}_{i}^{+}$, the likelihood change of its $t$-th action,
@@
 increases, where
-$\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}=\left\langle\mathbf{e}_{y^{+}_{i,t,k}}-\pi_{\theta(s)}(\cdot\mid x,y^{+}_{i,<t},o^{+}_{i,<t},y^{+}_{i,t,<k}),\mathbf{e}_{y^{-}_{j,t^{\prime},k^{\prime}}}-\pi_{\theta(s)}(\cdot\mid x,y^{-}_{j,<t^{\prime}},o^{-}_{j,<t^{\prime}},y^{-}_{j,t^{\prime},<k^{\prime}})\right\rangle$ and
+$\alpha^{-}_{(i,t,k),(j,t^{\prime},k^{\prime})}=\left\langle\mathbf{e}_{\bm{y}^{+}_{i,t,k}}-\pi_{\theta(s)}(\cdot\mid\bm{x},\bm{y}^{+}_{i,<t},\bm{o}^{+}_{i,<t},\bm{y}^{+}_{i,t,<k}),\mathbf{e}_{\bm{y}^{-}_{j,t^{\prime},k^{\prime}}}-\pi_{\theta(s)}(\cdot\mid\bm{x},\bm{y}^{-}_{j,<t^{\prime}},\bm{o}^{-}_{j,<t^{\prime}},\bm{y}^{-}_{j,t^{\prime},<k^{\prime}})\right\rangle$ and
 […]
@@
 is large. This establishes the trajectory-level statement in Theorem˜A.1.
+A.2 Formal Definition of the LLD Death Spiral
+Definition A.3 (LLD Death Spiral).
+[…]
+We define the token-level likelihood change at iteration $s$ as
+[…]
+and the trajectory-level likelihood change as
+$\Delta^{(s)}_{i}=\sum_{t=0}^{T}\sum_{k=1}^{|y^{+}_{i,t}|}\Delta^{(s)}_{i,t,k}$.
+The training dynamics are said to enter an LLD Death Spiral at iteration $s_{0}$ if there exist constants $\varepsilon<0$ and $\delta>0$ such that, for all $s\geq s_{0}$, the following conditions hold:
+  1. 1.
+    Persistent likelihood decay. The expected likelihood of correct trajectories decreases monotonically:
+    $\mathbb{E}_{i}\!\left[\Delta^{(s)}_{i}\right]\leq\varepsilon.$
+  2. 2.
+    Confidence erosion under response similarity. As likelihood decays, the model assigns no higher probability mass to incorrect trajectories,
+    $\mathbb{E}_{\hat{\bm{y}}^{-}\sim\pi_{\theta_{s}}(\cdot\mid\bm{x}_{i})}\!\left[\pi_{\theta_{s}}(\hat{\bm{y}}^{-}\mid\bm{x}_{i})\right]\;\leq\;\mathbb{E}_{\hat{\bm{y}}^{-}\sim\pi_{\theta_{s-1}}(\cdot\mid\bm{x}_{i})}\!\left[\pi_{\theta_{s-1}}(\hat{\bm{y}}^{-}\mid\bm{x}_{i})\right],$
+    while incorrect and correct trajectories remain representationally similar:
+    $\mathbb{E}_{\hat{\bm{y}}^{-},\hat{\bm{y}}^{+}\sim\pi_{\theta_{s}}(\cdot\mid\bm{x}_{i})}\!\left[\mathrm{sim}\!\left(\bm{h}(\bm{x}_{i},\hat{\bm{y}}^{-}),\,\bm{h}(\bm{x}_{i},\hat{\bm{y}}^{+})\right)\right]\geq c,$
+    for some constant $c>0$  [^1] In GRPO, correct and incorrect responses often exhibit similarity due to shared conditioning on the same query; see Figure 5..
+Here $\bm{h}(\cdot)$ denotes the policy’s hidden representation.
+  3. 3.
+    Self-amplifying decay. Lower response likelihood causes likelihood decay to become self-amplifying over training iterations.:
+    $\mathbb{E}_{i}\!\left[\Delta^{(s+1)}_{i}\right]\;<\;\mathbb{E}_{i}\!\left[\Delta^{(s)}_{i}\right$],
+    i.e., the magnitude of expected likelihood decay strictly increases over training.
+Lowlikelihood of Incorrect responses.
+We demonstrate the likelihood of correct and incorrect responses in Figure˜9(a).
+As shown, after step 80, the likelihood of incorrect responses is consistently lower than that of correct responses.
+This provides evidence that incorrect trajectories lie in a low-probability regime of the policy, which causes their token-level prediction errors to be assigned disproportionately large weights in the GRPO gradient.
+As a result, these low-likelihood incorrect responses exert an outsized negative influence on similar correct trajectories, suppressing the likelihood of correct actions and accelerating Lazy Likelihood Displacement.
+This observation is consistent with the theoretical analysis that low-probability negative samples amplify gradient interference and contribute directly to the onset of the LLD death spiral.
 Appendix B Additional Analysis
-B.1 Training Instability Prior to Collapse
-Figure 7: Demonstration of unstability prior collapse.
-While the final collapse is indeed triggered by large gradients from negative responses, substantial instability and a noticeable performance decline had already emerged earlier in training.
-As shown in Fig.˜7, the brown vertical line marks the first sharp drop in model reward (green curve), during which the likelihood of correct responses also decreases significantly, even though the gradient norm remains relatively small (around 2).
-As training progresses, the LD issue intensifies and the low confidence responses lead to gradient norm steadily grows, ultimately causing the model to collapse.
-These observations suggest that instability is present well before the collapse point and therefore warrants close attention.
-B.2 Impact of Regularization Strength
-To examine the effect of regularization strength, we ablate $\lambda\in{0,0.01,0.1}$ and plot the corresponding training reward dynamics in Fig.˜8.
-As shown, without regularization ($\lambda=0$), the model collapses around step 200.
-A small regularization value ($\lambda=0.01$, orange curve) delays but does not prevent collapse, which occurs around step 220.
-In contrast, a stronger regularization ($\lambda=0.1$) stabilizes training entirely, allowing the model to continue smoothly without collapse.
-We truncate the plot at step 250 for clarity, although training can proceed well beyond this point.
-Figure 8: Impact of regularization strength on Qwen2.5-7B-base.
-B.3 Unlocking Multi-step Reasoning via Answer Masking
-To further explore the potential of LLDS in promoting complex reasoning behaviors, we conducted an ablation study on the Qwen-2.5-3B-Base model, focusing on the regularization scope.
-Specifically, we mask out the final answer tokens from the calculation of the $L_{\text{LLDS}}$ term (denoted as "Mask Answer").
-It is important to note that Qwen-2.5-3B-Base inherently lacks the capability for multi-turn tool calling.
-As illustrated in Fig.˜9, this limitation is evident in the standard training setups: the baseline GRPO (blue curve) rapidly suffers from model collapse, with the number of valid searches dropping to zero.
-While the standard GRPO + LLDS (red curve) successfully stabilizes the training and prevents collapse, it fails to elicit any multi-step behavior, with the search frequency remaining stagnant at exactly 1.0.
-However, the "GRPO + LLDS-MA" variant (green curve) demonstrates a distinct emergent behavior.
-By removing the regularization constraint on the final answer tokens, we observe a significant increase in the number of valid searches per question, rising well above 2.0.
-This indicates that relaxing the penalty on the answer tokens effectively unlocks the model’s latent potential to engage in multi-step reasoning and utilize external tools—capabilities that standard methods failed to activate in the base model.
-Figure 9: Evolution of the number of valid searches per question on Qwen-2.5-3B-Base.
-The base model inherently lacks multi-turn tool calling capabilities.
-While the baseline (blue) collapses and standard LLDS (red) stabilizes at a single search step, LLDS-MA (green) successfully unlocks the model’s potential for multi-step reasoning, driving the number of valid searches significantly above 1.0.
-B.4 Qualitative Examples of Tool-Integrated Reasoning
-To further illustrate the behavioral characteristics of our trained models, we present qualitative reasoning traces generated by the Qwen2.5-7B family.
+B.1 Detail Related Work
+B.1.1 Tool-Integrated Reasoning and Agentic LLMs.
+Tool use has emerged as a powerful paradigm for equipping LLMs with adaptive reasoning capabilities.
+Early approaches relied on prompt-based orchestration (Lu et al., 2023; Shen et al., 2023) or multi-agent delegation frameworks to invoke tools without explicit training.
+Instruction-tuned models (Gou et al., 2023; Qin et al., 2023) later introduced structured tool-calling behaviors via supervised learning, but these systems remained largely static and were constrained to single-turn interactions.
+More recent work demonstrates that reinforcement learning can substantially enhance tool integration by enabling models to learn tool-usage policies through environmental feedback and task success.
+Notable systems such as RETool (Feng et al., 2025) and VERL-Tool (Jiang et al., 2025) support multi-step reasoning through dynamic tool use, self-verification, and error correction.
+This transition from static instruction-following to feedback-driven optimization has proven effective across a range of domains, including mathematical problem solving with code execution (Xue et al., 2025), open-domain question answering with retrieval (Jin et al., 2025), SQL generation from natural language (Jiang et al., 2025), and multi-modal visual reasoning (Gao et al., 2024).
+B.1.2 Training Collapse in Tool-Integrated GRPO.
+GRPO (Guo et al., 2025) has gained popularity in reinforcement learning due to its value-free, outcome-driven formulation.
+However, applying GRPO to train LLMs for multi-turn tool-integrated reasoning remains highly unstable (Jin et al., 2025).
+When initialized from base models using only verifiable rewards (Zero RL), GRPO-trained models often experience catastrophic failure, marked by abrupt reward collapse (Xue et al., 2025; Jin et al., 2025).
+In contrast, Proximal Policy Optimization (PPO) (Schulman et al., 2017) tends to exhibit more stable behavior under comparable settings (Jin et al., 2025).
+These collapse dynamics have been consistently observed across various systems, including Search-R1 (Jin et al., 2025), SimpleTIR (Xue et al., 2025), and ZeroSearch (Sun et al., 2025).
+While prior work has extensively documented these failures empirically, principled explanations remain limited.
+Among existing studies, (Xue et al., 2025) attributes the instability to low-likelihood incorrect responses that induce excessively large importance weights.
+Separately, (Liu et al., 2025) argues that off-policy accelerated inference introduces training–inference mismatch that exacerbates collapse.
+However, neither explanation accounts for why structurally similar algorithms such as PPO do not suffer comparable failures, nor do they clarify the deeper structural origin of these low-likelihood trajectories.
+Moreover, we observe that reward degradation often precedes the final collapse, suggesting a more fundamental optimization pathology rather than a purely stochastic failure.
+In this work, we identify LLD as the core mechanism underlying GRPO’s failure mode in multi-turn TIRL.
+To mitigate this issue, recent works introduce turn-level reward shaping for improved credit assignment (Zheng et al., 2025b; Zhang et al., 2025; Ji et al., 2025).
+StepSearch (Zheng et al., 2025b) constructs turn-wise rewards using ground-truth supporting documents, which are often unavailable in realistic settings.
+CriticSearch (Zhang et al., 2025) relies on an external LLM to judge each action, incurring additional computational cost and dependence on auxiliary model capability.
+Moreover, dense intermediate rewards could be susceptible to reward hacking (Guo et al., 2025).
+Recently, TreeGRPO (Ji et al., 2025), provide dense reward using tree-based rollout, but increases rollout complexity and computational overhead.
+In contrast, our approach shows that simple rule-based outcome rewards alone can achieve superior performance once GRPO training is properly stabilized.
+B.2 Additional Application Details
+B.2.1 Hyper-parameters
+Following Search-R1 (Jin et al., 2025), we implement our system using vLLM with a tensor parallelism degree of 1 and a GPU memory utilization ratio of 0.6.
+Rollout sampling is performed with temperature 1.0 and top-$p$ set to 1.0.
+We set the KL-divergence regularization coefficient to 0.001 and the clipping ratio to 0.2.
+For GRPO training, the policy LLM is optimized with a learning rate of $1\times 10^{-6}$, and five responses are sampled per prompt.
+Training is conducted on four H100 GPUs with a total batch size of 512, a mini-batch size of 256, and a micro-batch size of 64.
+The maximum sequence length is 4096 tokens, including up to 500 tokens for the generated response and 500 tokens for retrieved content.
+In addition, we reduce the maximum number of turns from 4 to 3 for efficiency.
+Although this change favors Search-R1, our method still outperforms it under this setting.
+During evaluation, the maximum action budget is set to 4, and the top-3 passages are retrieved by default.We set $\lambda=0.2$ for all models, except for Qwen-2.5-7B-Instruct where $\lambda=0.3$.
+For $\rm LLDS$-MA on Qwen-2.5-3B-Base, we linearly increase $\beta$ from 0 to $1$ according to $\beta=\min\!\left(\frac{V_{s}}{V_{B}},\,1\right),$ where $V_{s}$ denotes the average number of valid searches at time $s$ and $V_{B}$ is a baseline value fixed to $2$.
+B.2.2 Baseline Methods
+Direct Inference performs single-pass answer generation without external retrieval or iterative reasoning, serving as a minimal baseline to measure the intrinsic capability of the base LLM.
+Search-o1 (Li et al., 2025) introduces structured search planning with predefined search-action templates, enabling more controlled interaction with the retriever.
+However, it remains inference-only and does not learn to optimize search policies from task feedback.
+Retrieval-Augmented Generation (RAG) (Lewis et al., 2020) retrieves a fixed set of documents once at the beginning and conditions answer generation on the retrieved context.
+RAG improves factual grounding but lacks iterative refinement, making it less effective for multi-hop reasoning where intermediate queries must adapt to evolving reasoning states.
+R1 (Guo et al., 2025) is an RL-based fine-tuning method that optimizes reasoning trajectories using outcome-level rewards, but operates without a search engine.
+As a result, R1 primarily improves reasoning structure while remaining constrained by the model’s internal knowledge.
+Rejection Sampling (Ahn et al., 2024) generates multiple candidate reasoning trajectories and retains only those leading to correct answers for supervised fine-tuning.
+Search-R1-PPO / Search-R1-GRPO (Jin et al., 2025) extend R1 by incorporating a search engine during rollout and optimizing search-aware reasoning trajectories via PPO or GRPO.
+ZeroSearch (Sun et al., 2025) trains search-aware LLMs via reinforcement learning without interacting with real search engines, by replacing retrieval with a simulated LLM-based search model under a curriculum strategy.
+StepSearch (Zheng et al., 2025b) introduces turn-level reward shaping by assigning each search step an information-gain reward computed from similarity to ground-truth supporting documents, minus a redundancy penalty.
+This provides dense supervision but requires curated step-level document annotations that are often unavailable in realistic settings.
+We use the official checkpoints to evaluate.
+CriticSearch (Zhang et al., 2025) uses a frozen external LLM to retrospectively judge each search action as good or bad based on its contribution to the final answer, converting these judgments into turn-level rewards.
+While general and effective, it introduces additional computational cost, depends on the capability of the auxiliary judge model, and increases exposure to reward hacking.
+TreeGRPO (Ji et al., 2025) samples branching trajectories and propagates outcome rewards backward through the tree to produce step-level advantages via relative comparisons among sibling branches.
+This yields denser credit assignment but significantly increases rollout complexity and computational overhead.
+B.3 LLDS Variants
+LLDS: Token-Level Likelihood Preservation.
+One intial variant is the token-level penalty: within each preserved trajectory $\bm{y}_{i}$ and each action (turn) $\bm{y}_{i,t}=(y_{i,t,1},\dots,y_{i,t,|\bm{y}_{i,t}|})$, only likelihood-reducing tokens contribute to the loss.
+Then
+$L_{\rm LLDS(Token)}$ $=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{likelihood-reducing tokens}}.$ (9)
+This token-level selectivity ensures that the model is penalized only for genuinely harmful likelihood reductions, without interfering with updates that improve the response overall.
+However, some responses may still be globally improving even if a few individual tokens decrease; penalizing such cases may introduce overly strong constraints.
+LLDS: Response-Level Gating.
+To avoid penalizing globally improving actions, $\rm LLDS$ introduces a response-level gating mechanism: the penalty activates only when the total likelihood of the response decreases.
+The LLDS loss is:
+$L_{\rm LLDS(response)}$ $=\frac{1}{\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\sum_{t=0}^{T_{i}}|\bm{y}_{i,t}|}\sum_{\bm{y}_{i}\in\mathcal{Y}_{\rm pre}}\mathbbm{1}\!\left[\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\Delta_{i,t,k}>0\right]\sum_{t=0}^{T_{i}}\sum_{k=1}^{|\bm{y}_{i,t}|}\underbrace{\max(0,\Delta_{i,t,k})}_{\text{likelihood-reducing tokens}}.$ (10)
+However, a mismatch can arise between action-level and response-level likelihood changes: an individual action may suffer from LLD even when the overall response improves, or conversely, the response may exhibit LLD while only a subset of actions deteriorate.
+As a result, response-level gating can lead to either over-penalization or under-penalization at the action granularity.
+To address this issue, we adopt an action-level variant of $\rm LLDS$, defined in Equation˜4.
+Unless stated otherwise, $\rm LLDS$ refers to the action-level gating formulation throughout this paper.
+B.4 Additional Results
+B.4.1 Performance Comparison Among LLDS Variants
+We conduct an ablation study to compare different variants of $\rm LLDS$, including token-level, response-level, and action-level gating.
+The results are summarized in Table˜3.
+We first evaluate the three variants under the NQ-only training setting on Qwen2.5-3B-Base.
+The token-level variant yields the lowest performance among the three, indicating that penalizing all likelihood-reducing tokens, even within globally improving actions or responses, introduces overly strong regularization that interferes with policy optimization.
+We further compare the response-level and action-level variants under the more challenging NQ+Hotpot training setting on both Qwen2.5-3B-Base and Qwen2.5-3B-Instruct.
+In all cases, the action-level variant achieves the best performance across both general QA and multi-hop QA benchmarks, improving the overall average score by a consistent margin (e.g., from 0.353 to 0.360 on Qwen2.5-3B-Base and from 0.419 to 0.441 on Qwen2.5-3B-Instruct).
+While all variants outperform vanilla GRPO, action-level gating achieves the strongest empirical performance.
+Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
+NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
+Qwen2.5-3b-Base
+NQ-Only
+Search-R1-GRPO 0.440 0.582 0.413 0.478 0.265 0.244 0.061 0.113 0.171 0.303
++LLDS (token) 0.462 0.609 0.464 0.512 0.286 0.248 0.063 0.113 0.178 0.321
++LLDS (response) 0.478 0.605 0.449 0.511 0.288 0.250 0.062 0.129 0.182 0.323
++LLDS 0.491 0.615 0.450 0.519 0.306 0.275 0.067 0.121 0.192 0.332
+NQ+Hotpot
+Search-R1-GRPO 0.421 0.583 0.413 0.472 0.297 0.274 0.066 0.128 0.191 0.312
++LLDS (response) 0.480 0.631 0.455 0.522 0.352 0.311 0.087 0.153 0.226 0.353
++LLDS 0.479 0.627 0.453 0.520 0.345 0.323 0.082 0.210 0.240 0.360
+Qwen2.5-3b-Ins
+NQ-Only
++LLDS (token) 0.469 0.609 0.457 0.512 0.291 0.250 0.057 0.097 0.174 0.319
++LLDS (response) 0.490 0.610 0.450 0.517 0.294 0.193 0.074 0.121 0.171 0.319
++LLDS 0.485 0.614 0.455 0.517 0.299 0.266 0.062 0.129 0.189 0.330
+NQ+Hotpot
+Search-R1-GRPO 0.397 0.565 0.391 0.451 0.331 0.310 0.124 0.232 0.249 0.336
++LLDS (response) 0.462 0.622 0.460 0.515 0.432 0.383 0.180 0.395 0.347 0.419
++LLDS 0.462 0.627 0.468 0.519 0.443 0.447 0.197 0.444 0.383 0.441 (+31.3%)
+Search-R1-GSPO 0.376 0.561 0.374 0.437 0.332 0.329 0.098 0.250 0.252 0.331
++LLDS (response) 0.484 0.637 0.475 0.532 0.452 0.448 0.212 0.387 0.375 0.442(+33.5%)
++LLDS 0.489 0.640 0.483 0.537 0.466 0.456 0.215 0.403 0.385 0.451(+36.3%)
+Table 3: Results of different $\rm LLDS$ variants on General QA and Multi-Hop QA datasets.
+B.4.2 Results of training with NQ dataset only.
+In this section, we report results using training on the NQ dataset only.
+Models trained solely on NQ achieve strong performance on open-domain single-hop QA, but exhibit limited generalization to multi-hop reasoning.
+Incorporating HotpotQA (NQ+Hotpot) consistently enhances multi-step reasoning ability.
+As shown in Table˜3, for Qwen2.5-3B-Base, the GRPO baseline improves from 0.303 (NQ-only) to 0.312 (NQ+Hotpot), while $\rm LLDS$ (response) increases from 0.323 to 0.353, and $\rm LLDS$ further improves from 0.332 to 0.360 under the same setting.
+These gains indicate that adding multi-hop supervision from HotpotQA, where questions require multi-step reasoning to obtain correct rewards, encourages the model to retrieve, integrate, and reason over multiple pieces of evidence more effectively than training on NQ alone.
+B.4.3 Additional Results on Llama.
+In this section, we report results on Llama3.2-3B-Base.
+As shown in Table˜4, vanilla GRPO fails to learn tool use, collapsing into a “reason-and-answer” policy.
+This failure is reflected in its weak performance, achieving only 0.212 overall average and a particularly low 0.142 MH-Avg on multi-hop benchmarks.
+By incorporating $\rm LLDS$, the model is successfully stabilized and learns to invoke tools, yielding substantial gains across both general and multi-hop QA.
+Specifically, Gen-Avg improves from 0.305 to 0.532, MH-Avg rises from 0.142 to 0.255, and the overall average increases from 0.212 to 0.374.
+However, the learned tool behavior remains being single-turn usage.
+Finally, MA is applied on top of LLDS, the model is further encouraged to engage in multi-turn tool use, unlocking additional improvements on multi-hop reasoning.
+The MH-Avg increases from 0.255 to 0.343, while maintaining strong general QA performance (Gen-Avg = 0.538), leading to the best overall average of 0.427.
+Methods General QA Gen-Avg Multi-Hop QA MH-Avg Avg
+NQ TriviaQA PopQA HotpotQA 2Wiki Musique Bamboogle
+GRPO 0.233 0.461 0.220 0.305 0.196 0.232 0.044 0.097 0.142 0.212
+GRPO-LLDS 0.486 0.643 0.467 0.532 0.366 0.298 0.113 0.242 0.255 0.374
+GRPO-LLDS-MA 0.493 0.640 0.480 0.538 0.449 0.462 0.164 0.298 0.343 0.427
+Table 4: Results on General QA and Multi-Hop QA datasets for Llama3.2-3b-Base.
+(a) Likelihood of correct and incorrect responses.
+(b) Evolution of the number of valid searches per question on Qwen-2.5-3B-Base.
+Figure 9: Analysis of Likelihood dynamics and tool usage behaviors.
+Left: Likelihood comparison between correct and incorrect responses using NQ+Hotpot dataset.
+Right: Comparison of valid search steps across different training strategies.
+B.4.4 Encouraging Multi-step Reasoning via Answer Masking
+As shown in Figure˜9(b), the Qwen2.5-3B-Base model lacks inherent multi-turn tool-use capability: vanilla GRPO (blue) rapidly collapses, with search frequency fixed at 1.0, while GRPO+LLDS (red) stabilizes training but remains a single-turn search.
+In contrast, GRPO+LLDS-MA (green), which reduces regularization on answer tokens, exhibits clear emergent behavior, increasing valid searches per query beyond 2.0.
+This indicates that relaxing answer-token penalties unlocks latent multi-step reasoning and tool-use capabilities absent under standard training.
+B.5 Behavioral Characteristics
+To further illustrate the behavioral characteristics of our trained models, we present qualitative reasoning traces generated by the Qwen2.5-3B-Ins and Qwen2.5-7B-Ins.
 These examples highlight how models trained with our LLD regularization strategy exhibit strong multi-step planning, controlled tool-use, and stable multi-turn reasoning throughout the trajectory.
-Unlike standard GRPO models—which often suffer from premature collapse, our method enables the model to maintain coherent reasoning structure across search, verification, and final answer generation.
-As shown in Figures 10 and 11, the model not only performs step-by-step decomposition and self-verification but also executes consecutive search calls when needed, integrates retrieved evidence effectively, and produces a correct, concise final answer.
+Unlike standard GRPO models, which often suffer from premature collapse, our method enables the model to maintain coherent reasoning structure across search, verification, and final answer generation.
+As shown in  Figure˜10 and Figure˜11 , the model not only performs step-by-step decomposition and self-verification but also executes consecutive search calls when needed, integrates retrieved evidence effectively, and produces a correct, concise final answer.
 These qualitative behaviors align with our quantitative findings: stabilizing likelihood dynamics prevents the LLD Death Spiral, allowing the RL policy to leverage deeper tool-integrated reasoning without sacrificing robustness.
@@
 The only difference lies in the final answer token: the correct trajectory outputs the full entity “North Queensland Cowboys”, whereas the incorrect one truncates the name to “North Queensland”.
-Because GRPO assigns a single scalar reward to the entire trajectory, these two highly similar prefixes receive sharply different updates: the incorrect rollout pushes negative gradients onto tokens that are almost identical in embedding space to those of the correct rollout.
-This illustrates how small semantic deviations at the end of a trajectory can, through high prefix similarity, induce strong likelihood displacement on otherwise correct actions.
+Because GRPO assigns a single scalar reward to the entire trajectory, these two highly similar responses receive sharply different updates: The incorrect rollout pushes negative gradients onto tokens that are almost identical in embedding space to those of the correct rollout.
+This illustrates how small semantic deviations at the end of a trajectory can, through high similarity.
 Case 2: Low-likelihood, longer incorrect trajectories.
@@
 Because the likelihood of early tokens is extremely small, sampling drifts into semantically meaningless regions of the distribution, producing an extended sequence of nonsensical text.
-This also leads the model to violate the tool protocol, which subsequently triggers corrective feedback in the <information> channel.
+This behavior also causes the model to violate the tool protocol, triggering the system’s corrective rule in the <information> channel, without which the interaction fails.
 Although the model eventually issues a valid search, it still produces an incorrect answer (“The first-person narrator”).
@@
 Together, these examples concretely demonstrate how low-likelihood, overlong incorrect responses can dominate gradients and drive the LLD Death Spiral, even when the environment occasionally nudges the model back toward valid tool use.
-Question: Who was the screenwriter for Three Times? Ground Truth: Hou Hsiao-hsien SEARCH-R1: <think>
-To answer this question, I need to find out who the screenwriter for the movie Three Times is. I will conduct a search to gather this information.
-</think> <search> who was the screenwriter for Three Times </search> <information>
-Doc 1 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
-Doc 2 (Title: Paddy Chayefsky) Paddy Chayefsky Sidney Aaron "Paddy" Chayefsky (January 29, 1923 – August 1, 1981) was an American playwright
-Doc 3 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
+Appendix D More Discussion and Guideline
+Based on our analysis of Lazy Likelihood Displacement (LLD) and the emergence of the LLD Death Spiral, we consolidate several practical guidelines for stabilizing tool-integrated GRPO.
+Each recommendation follows directly from the core failure mechanisms identified in our study.
+Understand why tool-integrated GRPO is uniquely vulnerable to LLD.
+Unlike free-form RL on text, tool-augmented agents introduce structural conditions that magnify likelihood drift.
+First, tool calls inject inherently out-of-distribution tokens, such as search results, API outputs, or error messages, that differ sharply from pretrained language distributions.
+These OOD segments raise token-level uncertainty and make GRPO’s relative updates more volatile, accelerating the onset of likelihood displacement.
+Second, tool-based reasoning unfolds across multiple stages, with early stages (e.g., query formulation) stabilizing more quickly than later ones (e.g., tool-result interpretation).
+Because GRPO applies a single scalar reward to all tokens, early-stage tokens that share nearly identical prefixes across correct and incorrect trajectories receive conflicting gradient signals.
+This reward–token misalignment disproportionately harms stable prefixes and amplifies LLD.
+Recognizing these structural challenges is essential for diagnosing and mitigating instability in tool-integrated RL systems.
+Closely monitor likelihood dynamics:reward alone is insufficient.
+A central insight of our analysis is that likelihood degradation begins long before any visible drop in reward.
+Across models, reward continues to rise throughout both the early and steady-decay phases, even as the likelihood of correct responses is already drifting downward.
+Because GRPO updates are governed by likelihood ratios, such early degradation silently amplifies gradients and initiates instability.
+Monitoring only the reward therefore obscures the onset of LLD.
+In practice, reliable early-warning signals come from tracking action-level and trajectory-level log-likelihood, inspecting entropy trajectories, which surge sharply once likelihood begins to collapse, and watching for sudden spikes in likelihood ratios that indicate gradient amplification.
+Visualization of likelihood or entropy curves provides the earliest and most trustworthy indication that training is approaching the unstable regime.
+Considering Turn-Level Hidden Rewards as Future Work.
+Recent work by (Deng et al., 2026) introduces token-level hidden rewards to steer GRPO training toward exploration or exploitation.
+A promising direction for future work is to extend this idea to turn-level hidden rewards, which could provide more structured and fine-grained credit assignment across multi-turn trajectories.
+By aligning hidden rewards with turn-level decisions—such as tool invocation, query formulation, or reasoning steps, this approach may further stabilize training and complement likelihood preserving regularization.
+Question: What is the name of Kevin the teenager’s friend played by Kathy Burke in a series of sketches with Harry Enfield? Ground Truth: Perry SEARCH-R1: <think>
+To find the name of Kevin the teenager’s friend played by Kathy Burke in a series of sketches with Harry Enfield, I need to first identify the name of the series. Let’s start with that.
+</think> <search> Series of sketches with Harry Enfield and Kevin the teenager’s friend played by Kathy Burke </search> <information>
+Doc 1(Title: Kevin the Teenager) was ""Little Brother"", his father was played by Martyn Whitby and his mother was played by Caroline Quentin. Kevin’s mother (named Sheila in one of the earlier
+Doc 2(Title: Kevin the Teenager) Kevin the Teenager Kevin Patterson is a character created and played by the British comedian, Harry Enfield. Kevin originated in 1990 in the ""Little Brother""
+Doc 3(Title: Kevin & Perry Go Large) Farrelly Brothers’ humour."" Since the film’s release it has gone on to earn cult status amongst many clubbers across the UK and Ireland, largely for its
 </information>
 <think>
-From the search results, I found that the director of Three Times is Hou Hsiao-hsien.
-However, I need to find the screenwriter.
-Since the search results did not provide this information, I will conduct another search.
-</think> <search> screenwriter of Three Times </search> <information>
-Doc 1 (Title: Three Times) Three Times (Chinese: “Zuìhǎo de shíguāng”; lit. ‘Best of Times’) is a 2005 Taiwanese film directed by Hou Hsiao-hsien
-Doc 2 (Title: Eric Mendelsohn) Eric Mendelsohn (born November 1, 1964) is an American film director and screenwriter. Two of his films have been screened
-Doc 3 (Title: Paddy Chayefsky) are at the Wisconsin Historical Society and the New York Public Library for the Performing Arts, Billy Rose Theatre Division
+From the search results, it is clear that the series in question is "Harry Enfield’s Television Programme". Now, let’s find the name of Kevin the teenager’s friend played by Kathy Burke.
+</think> <search> Kevin the Teenager friend played by Kathy Burke in Harry Enfield’s Television Programme </search> <information>
+Doc 1(Title: Kevin the Teenager) boy named Perry Carter (played by the actress Kathy Burke, and based on an early character Burke portrayed on various Channel 4 shows). They starred in
+Doc 2(Title: Kevin the Teenager) was ""Little Brother"", his father was played by Martyn Whitby and his mother was played by Caroline Quentin. Kevin’s mother
+Doc 3(Title: Harry Enfield) Harry Enfield Henry Richard Enfield (born 30 May 1961) is an English comedian, actor, writer, and director. He is known in particular for his television
 </information>
 <think>
-From the search results, I found that the director of Three Times is Hou Hsiao-hsien.
-However, I need to find the screenwriter.
-Since the search results did not provide this information, I will infer that the screenwriter is Hou Hsiao-hsien, as he is the director of the film.
-</think> <answer> Hou Hsiao-hsien </answer>
-Figure 10: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search, with self-verification and finally answer the question.
-Question: Which was founded first University of California Santa Barbara or Hamdard University? Ground Truth: University of California, Santa Barbara SEARCH-R1: <think>
-To answer this question, I need to find out which university was founded first. I’$l$ conduct a search to gather this information.
-</think> <search> Which was founded first, University of California, Santa Barbara or Hamdard University </search> <information>
-Doc 1(Title: "University of California, Santa Barbara") University of California, Santa Barbara The University of California, Santa Barbara
-Doc 2(Title: "History of the University of California, Santa Barbara") History of the University of California, Santa Barbara The University of California, Santa Barbara (UCSB)
-Doc 3(Title: "History of the University of California, Santa Barbara") the chief executive was changed to Chancellor. UCSB’s first Provost was thus Clarence L.
+Based on the information provided, the name of Kevin the teenager’s friend played by Kathy Burke in Harry Enfield’s Television Programme is Perry.
+</think> <answer> Perry </answer>
+Figure 10: Qwen2.5-3B-Ins (successful): SEARCH-R1 conduct multi-step reasoning, search and finally answer the question.
+Question: What is the painful and incurable so far condition characterised by uterine cells growing outside the uterus? Ground Truth: Endometriosis SEARCH-R1: <think>
+I need to find out the painful and incurable condition characterized by uterine cells growing outside the uterus.
+</think> <search> painful and incurable condition characterised by uterine cells growing outside the uterus </search> <information> Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium, the layer of tissue that normally covers the inside of the uterus, grow
+Doc 2(Title: Endometrial cancer) Endometrial cancer Endometrial cancer is a cancer that arises from the endometrium (the lining of the uterus or womb). It is the result of the abnormal growth of cells
+Doc 3(Title: Uterine malformation) Uterine malformation A uterine malformation is a type of female genital malformation resulting from an abnormal development of the Müllerian duct(s)
+</information>
+<think> Based on the information provided, endometriosis seems to be the condition characterized by uterine cells growing outside the uterus.
+However, I need to confirm if it is painful and incurable.</think> <search> painful and incurable endometriosis </search> <information>
+Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium,
+Doc 2(Title: Endometriosis) A major symptom of endometriosis is recurring pelvic pain.
+The pain can range from mild to severe cramping or stabbing pain that occurs on both sides of the pelvis, in the lower back and rectal area, and even down the legs.
+Doc 3(Title: Endometriosis) to treat those whose symptoms are not manageable with other treatments.
+One estimate is that 10.8 million people are affected globally .
+Other sources estimate about 6–10% of women are affected.
+</information>
+<think> Based on the information, endometriosis is indeed a painful condition, and while there is no cure, treatments can help manage the symptoms.
+However, the question asks if it is incurable so far.
+I need to confirm if there is a known cure.
+</think> <search> incurable endometriosis </search> <information>
+Doc 1(Title: Endometriosis) Endometriosis Endometriosis is a condition in which cells similar to those in the endometrium,
+Doc 2(Title: Endometriosis) cases, girls may have endometriosis symptoms before they even reach menarche.
+The rate of recurrence of endometriosis is estimated to be 40-50% for adults over a 5-year period.
+The rate of recurrence has been shown to
+Doc 3(Title: Endometriosis) to treat those whose symptoms are not manageable with other treatments.
+One estimate is that 10.8 million people are affected globally .
+Other sources estimate about 6–10% of women are affected.
 </information>
 <think>
-From the search results, I found that the University of California, Santa Barbara traces its roots back to 1891. Now, I need to search for the founding date of Hamdard University.
-</think> <search> Hamdard University founding date </search> <information>
-Doc 1(Title: "Hamdard University") century. The university includes eight faculties, nine research institutes, three teaching hospitals and three affiliated engineering institutes.
-Doc 2(Title: "Hamdard University") Hamdard University Hamdard University (Urdu: ) () is a private research university with campuses in Karachi
-Doc 3(Title: "Hamdard University") electrical engineer Atta-ur-Rehman Memon, and Zillur Rahman (who has been serving as visiting professor of medicine since 1997)
-</information>
-<think>
-From the search results, I found that Hamdard University was founded in 1991.
-Now, I can compare the founding dates of the two universities.
-The University of California, Santa Barbara was founded in 1891, while Hamdard University was founded in 1991.
-Therefore, the University of California, Santa Barbara was founded first.
-</think> <answer> University of California, Santa Barbara </answer>
-Figure 11: Qwen2.5-7B (successful): SEARCH-R1 conduct multi-step reasoning, search and finally answer the question.
+Based on the information, endometriosis is not currently considered incurable, but treatments can help manage the symptoms. Therefore, the condition is not yet considered incurable.
+</think> <answer> endometriosis </answer>
+Figure 11: Qwen2.5-7B-Ins (successful): SEARCH-R1 performs multi-step reasoning and search with self-verification, identifies relevant documents (highlighted in bold), and then provides the final answer to the question.
 Question: Who won the NRL grand final in 2015? Ground Truth: North Queensland Cowboys SEARCH-R1:
@@
 Rightly summing up rudimentarily realizing computing Boolean cutely delayed indeed persuading predictable binding nutrients distinguishing vocabularily interrelational strategy cognizant broad range indicating integrating equally locating essence wholly sectioning open mainly umbrage confirming linguistically certainty personistically figuring hyperlinks meeting melodic benders incentives predictive spanning reunite nearest that strictly wholly consistent factual reformaturing calculating empirically traversing parameters meeting aesthetic
+</think>
 <information>